### v1

# SySLLM: Generating Synthesized Policy Summaries for Reinforcement Learning Agents Using Large Language Models

Abstract
Policies generated by Reinforcement Learning (RL) algorithms can be difficult to describe to users, as they result from the interplay between complex reward structures and neural network-based representations. This combination often leads to unpredictable behaviors, making policies challenging to analyze and posing significant obstacles to fostering human trust in real-world applications. Global policy summarization methods aim to describe agent behavior through a demonstration of actions in a subset of world-states. However, users can only watch a limited number of demonstrations, restricting their understanding of policies. Moreover, those methods overly rely on user interpretation, as they do not synthesize observations into coherent patterns. In this work, we present SySLLM (Synthesized Summary using LLMs), a novel method that employs synthesis summarization, utilizing large language models’ (LLMs) extensive world knowledge and ability to capture patterns, to generate textual summaries of policies. Specifically, an expert evaluation demonstrates that the proposed approach generates summaries that capture the main insights generated by experts while not resulting in significant hallucinations. Additionally, a user study shows that SySLLM summaries are preferred over demonstration-based policy summaries and match or surpass their performance in objective agent identification tasks.  [^1] All relevant resources for this study, such as SySLLM summaries, HIGHLIGHTS summaries, expert summaries, experience videos, the source code, and other materials are available at https://osf.io/c78jb/?view_only=76fbf59131d4409bac222e5929a654f2.
1 Introduction
Explainable Reinforcement Learning (XRL) methods aim to enable human users to understand agents’ policies and decision-making processes. Such understanding is crucial for fostering trust and ensuring the real-world deployment of RL agents. Several approaches to XRL have been developed in recent years [25], including saliency maps that visualize the agent’s attention, reward bars that break down expected utility into components, and causal explanations that clarify why actions were taken. Other methods extract logical rules to describe when actions are taken, but these approaches rely on predefined logical representations and often focus on individual actions rather than broader behavior patterns. Despite these advancements, experiments reveal that human users still struggle to form accurate mental models of agent behavior [6, 19, 23].
Recent advancements in large, pre-trained language models (LLMs) present a novel opportunity to enhance explanation methods for reinforcement learning (RL) policies. These models exhibit exceptional in-context learning abilities and possess extensive commonsense knowledge about human behavior, which makes them promising tools for generating meaningful, language-based explanations of RL policies. LLMs have already been employed in constructing world models and planning with domain-independent planners, showcasing their potential for reasoning and providing explanation [14].
Figure 1: Collecting the experience dataset (Section 4.2).
In this paper, we aim to leverage LLMs to generate comprehensive global explanations of RL agent policies. Specifically, our goal is to use LLMs to produce natural language summaries that encapsulate key patterns of behavior, capabilities, and limitations of RL policies. This approach seeks to build upon previous work that utilized demonstration-based policy summaries [4], where agent policies are described through selected trajectories in video. However, these methods are inherently constrained by the representativeness of the selected scenarios and the limited number of demonstrations that can feasibly be observed. They do not synthesize observations into coherent summaries, leaving users to rely heavily on their own interpretations to discern patterns. In contrast, we hypothesize that LLM-generated summaries can effectively distill these patterns into concise, coherent narratives, providing users with a clearer and more accessible understanding of agent policies.
Leveraging LLMs to generate accurate and comprehensible policy summaries is a non-trivial task. While LLMs exhibit remarkable reasoning abilities, they are prone to hallucinations—producing outputs that seem plausible but deviate from the agent’s true behavior [16]. This raises questions about whether such descriptions can reliably reflect the nuances of RL agent policies. Moreover, their utility heavily depends on end users’ ability to interpret them. Laypeople, who often lack familiarity with RL concepts, may struggle to form accurate mental models of agent behavior. Natural language explanations, while accessible, risk oversimplifying or failing to convey subtle patterns [6].
A further challenge lies in capturing the spatial and temporal relationships inherent in RL environments. These relationships evolve continuously over many steps and episodes: spatial dependencies include changing object positions, while temporal dependencies arise from sequential actions. Since LLMs are trained primarily on static text, they may struggle to model these dynamics effectively [24]. Together, these factors complicate the task of distilling and communicating agent behaviors in a way that remains faithful to the policy while being accessible to users.
Our method, Synthesized Summary using LLMs (SySLLM), leverages LLMs to generate comprehensive summaries of RL agent behavior. The proposed pipeline first collects the agent’s interactions within the environment (Figure 1) and converts these into natural language descriptions of observations and actions. These descriptions are then used to prompt an LLM, which synthesizes a natural language summary that captures key patterns in the agent’s behavior (Figure 2). We implement this methodology across five distinct environments, generating language-based summaries for agents with diverse behavioral patterns.
To evaluate SySLLM, we employ two complementary methods. First, we compare its summaries to those created by six RL experts, finding substantial agreement between the two. Second, we conduct a user study where participants interact with two types of policy summaries: visual summaries generated by the HIGHLIGHTS-DIV algorithm [3] and language-based summaries generated by SySLLM. Participants rated SySLLM higher in subjective preferences and performed equally well or better at identifying agents based on these summaries. These results highlight the effectiveness of SySLLM in both comprehensibility and practical utility.
The main contributions of our work are:
  • We introduce Synthesized Summary using LLMs (SySLLM), a novel and versatile method that generates holistic, synthesized, language-based summaries of RL agent policies using LLMs.
  • We validate the correctness of SySLLM by comparing its summaries to those of RL experts, showing substantial alignment and demonstrating that SySLLM accurately captures key policy details.
  • We conduct a large-scale user study, demonstrating that participants prefer SySLLM over a state-of-the-art demonstration-based policy summarization method in a summary preference task and perform equally well or better in an agent identification task.
Figure 2: Generating natural language global policy summaries (Section 4.3).
2 Background and Related Work
Explainable RL (XRL)
Our work extends the body of research concerned with explaining RL policies and particularly the line of work on global policy explanations (see Amitai and Amir [5] for a recent survey). Policy summarization approaches present users with demonstrations of the agent acting in the environment to convey its behavior. Several criteria for extracting such summaries have been proposed in the literature, including state importance, diversity, visitation frequency, and policy reconstruction [3, 17, 23, 18, 34]. While these approaches were shown to improve people’s ability to assess agent capabilities and predict agent behavior, users can only grasp a limited amount of demonstrations, and experiments show that their interpretation of policies highly depends on the subset of examples shown to them [23].
Other works propose learning a simplified model of the policy (e.g., a decision tree or a set of logical rules) that captures the essence of the agent’s decision-making [38, 8, 7]. Yet, these approaches are also limited due to the tradeoff between the accuracy of the simplified model and its interpretability to users.
We hypothesize that LLMs may offer a complementary approach to global explanations by capturing meaningful patterns of agent behavior and presenting them in understandable language to users.
Pre-trained Language Models in RL
Large Language Models (LLMs) are probabilistic text models trained on substantial text datasets. Models like GPT [30], which are autoregressive, excel in text generation by sequentially predicting future words based on the preceding text. Such models can be useful in the context of RL as they encode common-sense knowledge about the world. Conversely, models such as BERT [12], which are based on the masked framework, focus on understanding deep sentence semantics by training to recognize parts of text that are intentionally obscured. Encoder-decoder models merge these traits to adeptly handle tasks requiring both the interpretation and creation of text. These models have been extensively pre-trained on large-scale text corpora, equipping them with the ability to tackle complex language processing tasks with minimal specific instruction, often achieving impressive results [9]. This extensive training allows them to efficiently process and summarize complex narratives, and to identify relevant patterns within large datasets.
Recent work has developed methods that use LLMs in the training of RL policies. For example, LLMs are being used to guide intrinsic exploration by suggesting plausibly useful goals and providing intrinsic rewards [13, 27]. Other works use LLMs to generate useful state abstractions [29, 28]. Additionally, LLMs are employed to provide natural language instructions and dynamic feedback, enhancing agent learning in complex, open-ended environments [39]. Differently from these approaches, in this work, we aim to use LLMs to explain RL policies.
3 Problem Statement
We define a partially observed Markov Decision Process (POMDP) as a tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{O},T,O,R,\gamma\rangle$, where observations $o\in\mathcal{O}$ are derived from states $s\in\mathcal{S}$ and actions $a\in\mathcal{A}$ via $O(o|s,a)$. The dynamics are governed by the transition function $T(s^{\prime}|s,a)$, with $R$ as the reward function and $\gamma$ as the discount factor. POMDPs generalize Markov Decision Processes (MDPs), where observations fully reveal the state.
The agent operates under a stochastic policy $\pi:\mathcal{O}\rightarrow\Delta(\mathcal{A})$, mapping each observation $o\in\mathcal{O}$ to a probability distribution over actions $\mathcal{A}$. The objective of the policy is to maximize the expected cumulative discounted reward.
Given a POMDP environment and the agent’s policy, our goal is to generate a natural language summary that encapsulates the agent’s strategy and decision-making process. This summary explains the policy’s operational dynamics, effectiveness, and behavior across scenarios. Formally, a summarization is a mapping from $(\mathcal{M},\Pi)$ to $\mathcal{T}$, where $(\mathcal{M},\Pi)$ represents environment-policy pairs, and $\mathcal{T}$ is the set of natural language summaries.
4 Methodology
Our method consists of several key components, which we formalize in Algorithm 1. In the first phase (illustrated in Figure 1), we run the agent in its environment and perform a captioning procedure to convert experiences into natural language. These captioned experiences are logged into an experience dataset (Sections 4.1 and 4.2). In the second phase, which we illustrate in Figure 2, we take the experience dataset as input and output a natural language summary that aims to capture patterns in the agent’s policy (Section 4.3). To do so, we compose a prompt with the formatted experience dataset that is then sent to an LLM to generate the summary.
4.1 Captioner
As will be detailed in Section 4.2, our approach utilizes captioners to leverage LLMs for summarizing an agent’s policy. We convert each observation $o\in\mathcal{O}$ and action $a\in\mathcal{A}$ into natural language descriptions that highlight their key features. In this subsection, we formalize the role and function of the captioners.
Observation Captioner:
The observation captioner, $C_{\text{observation}}$, translates observations into natural language descriptions:
$C_{\text{observation}}:\mathcal{O}\rightarrow\Sigma^{\star}.$
Where $\Sigma^{\star}$ represents the set of all possible strings. This function distills complex state information into comprehensible text, such as translating a grid map into a description of object locations.
Action Captioner:
Similarly, the action captioner, $C_{\text{action}}$, maps actions to natural language representations:
$C_{\text{action}}:\mathcal{A}\rightarrow\Sigma^{\star}.$
Similar to past work [26, 20] we assume the availability of a captioner language annotator. The process of generating captions is supported by a range of established methods, which are further discussed in Section 8.
4.2 Collecting the Agent’s Experience Dataset
We describe the first phase of Algorithm 1, detailed in Lines 114.
We assume that the environment ($ENV$) supports two functions. The first function is RESET, which resets the environment to its initial state and returns the initial observation, and STEP, which takes an action as input and returns the next observation, the reward for that action, a boolean indicating whether the episode has ended, and additional information.
First, we initialize the experience dataset $ED$ (Line 1). The for loop (Line 2) iterates over multiple episodes, initializing the step counter $t=1$ at the beginning of each episode (Line 3) and obtaining the initial observation $o_{1}$ (Line 4). The variables $epReward$ and $done$ are also initialized to track the cumulative reward and episode completion status, respectively (Lines 56). The while loop (Line 7) continues until the episode concludes.
During each iteration of the while loop, an action is sampled from the policy $\pi$ (Line 8), and both the observation and action are converted into natural language representations by the captioners. These representations are added to the experience dataset along with the cumulative reward (Line 9). Subsequently, the environment is updated using the selected action, and the next observation is retrieved (Line 10).
Finally, the cumulative reward and step index are updated accordingly (Lines 1112). The information stored in the experience dataset is detailed in Section B of the Appendix.
4.3 Generating Natural Language Summaries
We move on to addressing the second phase of Algorithm 1, covered in Lines 1517.
To generate natural language summaries, we use the experience dataset and an LLM. The process begins by formatting the experience dataset into a structured format suitable for integration into a prompt. This format captures the sequential and contextual details of the agent’s interactions, creating a comprehensive narrative of its decision-making process (Line 15).
The prompt design, inspired by Chain-of-Thought reasoning [40], decomposes complex behavioral analysis into specific components. By incorporating detailed environmental parameters and quantitative evaluation guidelines, the prompt effectively guides the model’s analysis, enhancing its ability to generate insightful summaries of reinforcement learning behavior. This design was refined through iterative exploration, as detailed in Section H of the Appendix.
The prompt structure consists of four main components:
  • General Instructions: for example, “Generate a summary that captures the essence of the agent’s policy.”
  • Environment Description: such as, “Describe the agent’s task and key aspects of the environment.”
  • Interpretation Instructions: e.g., “Descriptions are from the agent’s perspective.”
  • Output Instructions: for instance, “Provide a concise summary.”
The LLM then produces the summary based on the prompt and the formatted experience dataset (Line 16).
Algorithm 1 SySLLM algorithm
Input: Environment $ENV$, trained policy $\pi$, captioners $C_{\text{observation}}$, $C_{\text{action}}$
Parameter: number of episodes $N$, prompt $P$
Output: summary $\in\mathcal{T}$
1:  Initialize experience dataset $ED$
2:  for $i=1$ to $N$ do
3:     Let $t\leftarrow 1$
4:     $o_{t}\leftarrow$ $ENV$.RESET()
5:     $epReward\leftarrow 0$
6:     $done\leftarrow\text{False}$
7:     while $\neg done$ do
8:        $a_{t}\sim\pi(\cdot\mid o_{t})$
9:        $ED$.ADD($C_{\text{observation}}(o_{t})$, $C_{\text{action}}(a_{t})$, $epReward$)
10:        $o_{t+1},r_{t+1},done\leftarrow$ $ENV$.STEP($a_{t}$)
11:        $epReward\leftarrow epReward+r_{t+1}$
12:        Let $t\leftarrow t+1$
13:     end while
14:  end for
15:  $formattedInput$leftarrow $P$ + FORMAT($ED$)
16:  $summary$leftarrow LLM($formattedInput$)
17:  return $summary$
5 Implementation
We demonstrate our approach in five diverse environments to evaluate summaries for seven agents, each with a unique policy. Specifically, we use the MiniGrid framework [10].
5.1 Captioners
To convert raw observations and actions into natural language, we designed a rule-based captioning system for the agent’s grid-based representation of the environment. This system systematically generates detailed descriptions of visible elements, emphasizing spatial relationships for navigation. Each action in the environment’s discrete action space is translated into natural language annotations, such as ‘move forward’ or ‘turn left’. We evaluated the effectiveness of this captioner by experimenting with and refining its design, ensuring the generated descriptions are accurate and comprehensive.
5.2 Agents’ Policy Training
We trained three qualitatively different agents for the MiniGrid-Unlock environment:
  • Goal-directed Agent: Follows a policy designed to minimize the number of steps to unlock the door efficiently. The agent can see all grid cells within a 180-degree field of vision in front of it.
  • Short-Sighted Agent: Follows a policy designed to minimize the number of steps to unlock the door efficiently but the agent’s sight is $3\times 3$.
  • Random Agent: Selects actions uniformly from the action space. We chose this policy to see how the LLM summarization process handles cases where there is no clear pattern of behavior.
Additionally, we trained four other agents in the following MiniGrid environments: Dynamic Obstacles, Lava Gap, Red-Blue Doors, and Crossing. Using SySLLM, we generated a policy summary for each of these agents. All of the agents are trained using the PPO algorithm [33, 31]. An overview of agent performance is shown in Table 1. The mean reward for the first three agents in the Unlock environment shows a declining trend, which aligns with the policies that each of these agents employ. In addition, the success rate of the agents in the other environments remains consistently high. Implementation details are provided in Section E of the Appendix.
Agent Mean Reward ± SD Mean Length Success Rate
Unlock Goal-directed 0.73 ± 0.21 20.25 0.93
Unlock Short-sighted 0.41 ± 0.27 44.43 0.77
Unlock Random 0.00 ± 0.01 70.00 0.00
Dynamic Obstacles 0.78 ± 0.06 17.20 1.00
Lava Gap 0.896 ± 0.02 10.82 1.00
Red Blue Doors 0.70 ± 0.26 17.06 0.88
Crossing 0.67 ± 0.18 24.80 0.94
Table 1: Report of agents’ performance metrics for 500 episodes.
5.3 Generating Policy Summaries
For each agent, we collect an experience dataset that logs the captioned observations, actions taken by the agent, and additional relevant data. Specifically, we log 50 episodes for each agent. After collecting and formatting the data, we synthesized the narratives into coherent global policy summaries for each agent using zero-shot prompting. We employed the gpt-4-turbo model with a temperature of $0.2$ [1]. The prompt used is shown in Section C of the Appendix.
In selecting the final summary, we employed a methodical approach by first generating ten distinct summaries through Line 16 of our algorithm. Each of these summaries was then embedded into a high-dimensional vector space using text-embedding-3-small, allowing us to quantify the semantic proximity of each summary to the others within this space. The centroid of these embeddings was calculated to identify the geometric mean of the summaries, representing the central tendency of the semantic dimensions explored by the model.
We chose the summary that is the median (5th) closest to the centroid. This position represents a solution to the trade-off between generality and specificity. Summaries closer to the centroid tend to include more general statements that apply broadly across different scenarios, while those further away incorporate more specific details unique to particular contexts. An example of a selected summary is provided in Section D of the Appendix.
5.4 Insights From Summaries
Using the SySLLM method, we generated summaries capturing key aspects of each agent’s decision-making and behavior. This section presents key insights from these summaries, concisely illustrating agents’ policies and critical action patterns. Table 2 provides a overview of these insights, showcasing their alignment with task objectives and how agents behave across various environments.
For the goal-directed agent in the Unlock environment, Insight (2) states that the agent “demonstrates a pattern of turning towards the nearest key or door once identified, suggesting a straightforward heuristic of minimizing distance to the target.” This reflects the agent’s efficient and consistent behavior, dynamically adjusting its path to achieve objectives with minimal steps. Insight (3) quantifies this efficiency, noting that episodes are completed within 15-25 steps on average, often achieving near-maximum rewards. These observations are validated by performance metrics in Table 1, which report a mean episode length of 20.25 and a high mean reward of 0.71. Together, these results highlight SySLLM’s ability to capture both qualitative patterns and quantitative performance metrics.
In the Dynamic Obstacles environment, SySLLM captures the agent’s ability to navigate challenges effectively. Insight (1) states that the agent “consistently demonstrates the ability to avoid obstacles (blue balls) by making turns or moving forward when the path is clear.” This highlights its proficiency in avoiding collisions, a critical success factor. However, Insight (3) notes inconsistent behavior when navigating closely spaced obstacles, with the agent occasionally taking longer routes or making unnecessary turns. These observations demonstrate SySLLM’s ability to provide a nuanced understanding of both the agent’s strengths in straightforward scenarios and its struggles in more complex configurations.
Agent Insights from Summaries
Unlock
Goal-Directed
(1) The agent effectively identifies keys and adjusts its path based on their relative position, shifting focus to unlocking the door.
(2) It consistently turns towards the nearest key or door, minimizing distance, which remains consistent across episodes.
(3) The agent completes episodes efficiently, averaging 15-25 steps with near-maximal cumulative rewards.
Unlock
Short-sighted
(1) The agent follows a right-wall method, moving forward until encountering an obstacle before turning.
(2) It identifies keys and doors efficiently, maneuvering toward and using them correctly.
(3) Decisions are heavily influenced by its immediate field of vision, reacting only to nearby objects.
Unlock Random (1) The agent exhibits unstructured behavior, often repeating unnecessary actions.
(2) It frequently toggles doors multiple times or picks up and drops keys without using them effectively.
Lava Gap (1) The agent consistently avoids lava, demonstrating awareness of environmental hazards.
(2) Upon encountering an obstacle, it either turns or moves in the opposite direction.
Red-Blue Doors (1) The agent prioritizes opening the red door before the blue door, optimizing reward accumulation.
(2) It successfully interacts with doors in a structured sequence, adhering to task constraints.
Crossing (1) The agent moves towards the green goal once it enters its field of vision, adjusting its path accordingly.
(2) It avoids collisions with walls through timely directional changes.
Dynamic
Obstacles
(1) The agent effectively avoids moving obstacles (blue balls) by adjusting its movement.
(2) It identifies objects in its field of vision and makes informed navigation decisions.
(3) In dense obstacle scenarios, occasional inefficiencies or unnecessary turns are observed.
Table 2: Summarized insights on agent behaviors across different tasks.
6 Expert Evaluation of Summaries
To validate the quality and alignment of SySLLM summaries, we conducted evaluations with RL experts. Specifically, six graduate students actively engaged in RL research with experience in training agents were recruited. These experts were divided into two groups: Experts 1, 2, and 3 evaluated the Minigrid-Unlock goal-directed and short-sighted agents, while Experts 4, 5, and 6 evaluated the remaining agents. Each expert was shown videos approximately 120 seconds long, depicting the agents’ behaviors in the environment. Based on these videos, the experts were tasked with summarizing the agents’ behaviors. To ensure a fair comparison with SySLLM-generated summaries, we provided the experts with instructions that aligned closely with the prompts used for the language model (see Section F of the Appendix for detailed instructions).
To evaluate how much of the experts’ observations were captured by the SySLLM summaries, we used a recall metric. This metric quantified the extent to which the SySLLM summary covered the points identified by the experts. Each expert summary and the SySLLM summary were broken down into lists of key points, and a scoring system was applied: 1 point for a match with a SySLLM key point, 0.5 points for a partial match, and 0 points for no match. The recall scores, as shown in Table 3, ranged from 0.687 for the Unlock goal-directed agent, indicating moderate agreement, to 0.914 for the Crossing agent, reflecting a near-complete capture of expert insights. The overall mean recall score across all agents was 0.814, demonstrating substantial coverage of the experts’ points by the SySLLM summaries.
In the second evaluation phase, the experts were presented with key points from the SySLLM summaries that they had not mentioned in their own summaries. They were tasked with classifying these points as Matched, Partially Matched, or Not Matched based on their agreement with the content. Using the same scoring system as for recall (Matched = 1, Partially Matched = 0.5, Not Matched = 0), we calculated a precision score for each summary. The precision score measures how many of the SySLLM summary points were validated as correct by the experts. Importantly, the precision calculation includes all SySLLM key points, encompassing both the points classified in this phase and those already evaluated during the recall assessment. The precision scores, as shown in Table 3, ranged from 0.769 (Dynamic Obstacles agent) to 0.864 (Short-Sighted agent), with an overall mean precision score of 0.819 across all agents. In addition to quantifying correctness, this metric offers a way to evaluate potential hallucinations by identifying points that may not align with expert observations. The high precision scores suggest that hallucinations were minimal in our evaluations. While instances of hallucinated content in SySLLM summaries were rare, they underscore the inherent challenges in summarizing RL agent behaviors using LLMs. One example from the Crossing agent summary: “The agent frequently checks for walls in its path and adjacent tiles.”. This point is classified as “not matched” by all experts, indicating a possible hallucination.
Taken together, the integration of recall and precision metrics demonstrates SySLLM’s capability to produce comprehensive and accurate summaries that closely align with expert observations.
Agent Expert Recall Precision Mean Recall Mean Precision
Unlock Goal-Directed 1 0.500 0.864 0.687 0.864
2 0.643 0.864
3 0.917 0.864
Unlock Short-Sighted 1 0.800 0.846 0.878 0.839
2 0.833 0.807
3 1.000 0.923
Dynamic Obstacles 4 0.583 0.692 0.739 0.769
5 0.833 0.692
6 0.800 0.923
Lava Gap 4 0.667 0.769 0.794 0.811
5 0.786 0.846
6 0.929 0.818
Red Blue Doors 4 0.857 0.767 0.871 0.834
5 0.857 0.867
6 0.900 0.867
Crossing 4 0.750 0.731 0.914 0.795
5 0.917 0.808
6 1.000 0.846
Table 3: Recall and Precision of Expert Insights for SySLLM.
7 User Study
To evaluate the usefulness of the SySLLM summaries, we conducted a user study comparing SySLLM summaries to HIGHLIGHTS-DIV (hereafter referred to as “HIGHLIGHTS”) summaries, which is a standard video-based benchmark in XRL [3]. For the study, we focused on the three agents from the MiniGrid Unlock environment: goal-directed agent, short-sighted agent, and random agent. These agents were chosen to represent qualitatively distinct behaviors, ensuring a diverse set of policies for participants to evaluate. This variety allowed us to better assess how well the summaries captured different decision-making strategies and provided interpretable insights into the agents’ behavior. For completeness, we provide an intuitive explanation of HIGHLIGHTS in Section A of the Appendix. We assessed participants’ subjective preferences as well as their ability to identify agent behavior based on the summaries.
7.1 Procedure
Participants were first shown a tutorial explaining the rules of the Unlock environment. They then had to pass a quiz ensuring they read and understood the rules. Next, they were asked to complete two different tasks.
Task 1: Summary Preferences
In the first task, participants evaluated and compared summaries of agents based on their policies. Initially, they watched a 120-second video showcasing full episodes to familiarize themselves with the agent’s policy. Following this, participants were presented with the first summary, either from the SySLLM or HIGHLIGHTS condition (see Section 7.2), and rated their agreement on a series of metrics using a 7-point Likert scale. The metrics, adapted from Hoffman’s explanation satisfaction questionnaire [15], included Clarity, Understandability, Completeness, Satisfaction, Usefulness, Accuracy, Improvement, and Preference. These adaptations were tailored to the study’s specific needs and excluded irrelevant items (see Section F for the full scale).
Participants then evaluated the second summary, corresponding to the alternative condition (HIGHLIGHTS or SySLLM), using the same set of questions. This allowed for a direct comparison between the two summary types. After assessing both summaries, participants answered additional questions that compared the summaries and expressed their preferences, highlighting which summary they found more helpful and explaining their reasons.
Task 2: Agent Identification
While the first task evaluated participants’ subjective perceptions of the summaries, the second task assessed their objective ability to match a given summary to the correct agent. Participants were presented with summaries, each linked to either the SySLLM or HIGHLIGHTS condition, and shown three short experience videos (approximately 20 seconds each) depicting different agents’ actions in the environment. They were asked to identify which video corresponded to the behavior described in the summary. Participants indicated their chosen video and rated their confidence on a 1-7 Likert scale, along with justifications for their decision in an open text format.
To keep the experiment concise and focused on key comparisons, we set up three matching questions—labeled $Q1$, $Q2$, and $Q3$. Each question featured a summary of a different agent: the goal-directed agent in $Q1$, the random agent in $Q2$, and the short-sighted agent in $Q3$. Depending on the condition to which a participant was assigned, they encountered either a text-based summary (SySLLM) or a video summary (HIGHLIGHTS) for these agents. In each question, alongside the video of the associated agent, participants were also shown two additional videos featuring different, unrelated agents. The task for the participants was to correctly identify the agent that matched the given summary. Success was measured by the participant’s ability to accurately associate the correct agent with its respective summary.
7.2 Experimental Conditions
Participants were divided into four conditions, varying by summary type (SySLLM or HIGHLIGHTS) and agent type (goal-directed or short-sighted). For Task 1, we used a within-subject design where participants viewed summaries of an agent’s behavior in two sequences. To counter ordering effects, one group received the SySLLM summary first and the HIGHLIGHTS summary second, while the other group received the reverse order. In Task 2, we employed a between-subject design. Participants viewed only SySLLM or HIGHLIGHTS summaries, corresponding to the sequence they were exposed to in Task 1. This setup ensured that participants experienced one of four unique combinations of summary and agent types in a controlled manner. Details of these conditions are shown in Table 4.
Benchmark
For generating the HIGHLIGHTS summaries, we used the following parameters: 300 traces (episodes), a context length of 5, and a total of 20 highlights to effectively capture key behaviors and the necessary context within the Unlock environment. These choices ensure that the summaries provide a detailed yet concise depiction of critical actions, appropriately matching the environment’s scale and complexity.
Condition Task 1 Sequence Task 1 Agent Type Task 2 Summary Type
1 SySLLM $\rightarrow$ HIGHLIGHTS Goal-directed Agent SySLLM
2 SySLLM $\rightarrow$ HIGHLIGHTS Short-sighted Agent SySLLM
3 HIGHLIGHTS $\rightarrow$ SySLLM Goal-directed Agent HIGHLIGHTS
4 HIGHLIGHTS $\rightarrow$ SySLLM Short-sighted Agent HIGHLIGHTS
Table 4: Experimental conditions.
7.3 Participants
We recruited participants through Prolific (N = 200). Participants were native English speakers from the US, UK, Australia, or Canada. Participants received a 3.75 pound base payment and an additional, 1 pound bonus if participants answered all questions correctly in Task 2. Participants who failed the attention question were excluded from the final analysis, as well as participants who completed the survey in less time than the length of the videos presented. After exclusions, we had 192 participants (94 female, Mean age = 36.41, STD = 12.05).
7.4 Results
Task 1: Summary Preferences
The SySLLM summaries consistently received higher ratings compared to the HIGHLIGHTS summaries across all metrics (see Figure 3). To assess the statistical significance of the differences in ratings, we first calculate the average scores for each participant under both the SySLLM and HIGHLIGHTS conditions. We then compute the difference between these averages for each participant. A paired t-test is conducted on these differences to evaluate whether the SySLLM condition significantly outperforms the HIGHLIGHTS condition across all metrics. The results indicate a statistically significant difference in favor of the SySLLM condition, with statistic $T=13.99$ and $p<10^{-33}$.
Following the evaluations of specific summaries, we asked the participants to subjectively compare the textual and video summaries. Specifically, we asked them which summary they think is better for understanding the agent’s overall behavior based on the observed video of a large sample of trajectories. They were further asked to rate the extent to which they preferred the textual (or visual) summary (on a 1-7 Likert scale). The results we obtained are conclusive. The responses indicate a clear preference for the SySLLM summary, with 75.5% of participants favoring the language-based summaries. Regarding the second question, the mean rating was 5.97 with a standard deviation of 1.44. In the 1-7 Likert scale, a score of 4 represents indifference to the different policies; hence, a mean score of 5.97 indicates a substantial preference for the textual summary.
Qualitative Analysis The open text feedback we received from participants of Task 1 turned out to be highly valuable. Some participants explained that while HIGHLIGHTS summaries required their own interpretation of the videos, SySLLM summaries provided an interpretation of the agent’s policy, e.g., “the [SySLLM] summary was thorough in explaining the movement of the agent. It describes information that isn’t visible in the video. The information in the summary was objective while watching the video summary and interpreting it is based on my view which is subjective.” Other participants mentioned that SySLLM summaries also explained why the agent acted as it did. For instance, one participant noted that “there are instances in the video where the agent seems to turn random corners instead of straight as you would expect it to. The summary explains why.” Another participant wrote that “the video itself does not fully explain but rather how you interpret what it is doing and in the case of errors it starts doing something wrong you don’t know what causes it”.
Additionally, some participants liked that SySLLM summaries provided more context, e.g., “the textual summary provides context for what is occurring. Conversely, the video summary doesn’t inform the viewer of the ’behind the scenes’ that is occurring. From the textual information, I identified many strategies and code that the agent is using such as using the maze strategy to always turn right when it hits a wall. On the other hand, the video does not provide this information unless you have prior knowledge such as coding experience on conditionals and iterations to know what it occurring.”
Figure 3: Participant scores for Task 1. Each metric includes a box plot of the Likert scores. Notably, SySLLM receives higher ratings than HIGHLIGHTS across all metrics.
Task 2: Agent Identification
As shown in Figure 4, participants’ ability to correctly identify the agent based on the provided summaries varied between the SySLLM and HIGHLIGHTS conditions across the three questions.
For $Q1$ ($\chi^{2}=1.0523$, $p=0.3050$, odds ratio (OR) = 0.5843), $Q2$ ($\chi^{2}=0.6703$, $p=0.4129$, OR = 0.7550), and $Q3$ ($\chi^{2}=0.7279$, $p=0.3936$, OR = 1.3875), the Chi-Square tests showed no statistically significant differences. Independent t-tests showed no statistically significant differences in confidence scores between the SySLLM and HIGHLIGHTS conditions for $Q1$ ($t=-0.6202$, $p=0.5359$) and $Q2$ ($t=0.5565$, $p=0.5785$). However, for $Q3$, there was a statistically significant difference ($t=3.4197$, $p=0.0008$).
To conclude, participants did not show a significant preference for the SySLLM or HIGHLIGHTS condition when identifying the correct agent based on the summaries provided, except for $Q3$ where the SySLLM condition had significantly higher confidence scores. Further investigation may be required to understand the nuances behind these results and whether different summary methods or conditions might yield different outcomes.
Figure 4: Correctness rate in Task 2. For each question, we measure the proportion of users that identified the correct agent under both SySLLM, HIGHLIGHTS, and a random guess. For each bar, we included the 95% confidence intervals. Notice that both SySLLM and HIGHLIGHTS receive higher scores than the random guess. For all three questions, the difference in correctness rate between SySLLM and HIGHLIGHTS is not statistically significant.
8 Discussion and Future Work
We introduce SySLLM, leveraging LLMs to generate summaries of reinforcement learning agents. Its effectiveness is demonstrated across five distinct environments, including seven different agents. Expert evaluations show that SySLLM closely aligns with expert-generated summaries, capturing key aspects of agent behavior. User studies reveal that participants rate SySLLM summaries higher than HIGHLIGHTS on subjective metrics and perform equally well in agent identification tasks. These results provide evidence for the potential usefulness of the proposed approach. We now turn to discuss several limitations of the SySLLM method and possible ways to address them, as well as additional directions for future work.
SySLLM relies on the availability of a captioning function, a widely accepted approach in language-integrated methodologies, as evidenced by prior research [13, 26, 20]. In simulated environments, captioning is particularly straightforward, as ground truth states are readily accessible, and numerous RL environments, such as NetHack/MiniHack [32, 22] and text-based games [35, 11], already incorporate language-based features. In more complex or real-world scenarios, robust techniques such as object detection [41], cutting-edge captioning frameworks [36], action recognition systems [21], or advanced vision-language models [2] can be utilized. Alternatively, annotator models [13] or pretrained foundation models [37] offer viable solutions for generating captions. With continued advancements in general-purpose captioning models, we anticipate that readily available, off-the-shelf captioning tools will soon support an even broader range of tasks, further solidifying the practicality of this approach.
Advancements in LLMs, particularly larger context windows, can extend this approach to large-scale and more complex environments, further validating its effectiveness and unlocking its potential to provide deeper insights into agent behavior.
Future work could explore a question-answering (QA) framework, allowing users to query the agent’s policy directly. For example, users could ask why specific actions were taken or predict the agent’s behavior in new scenarios. This interactive system would provide deeper insights into the agent’s decision-making process, offering a flexible and intuitive way to understand its behavior.
Another avenue for future work is generating multimodal summaries, combining natural language explanations with visual representations of key moments, trajectories, and decisions. That is, for each pattern identified by the LLM, the summary would also include demonstrations of the pattern extracted from agent traces. This approach could enhance user understanding by integrating text and visuals to more effectively convey complex policies. Visual elements from the agent’s experience dataset might help mitigate hallucinations by grounding explanations in observable behavior, potentially improving reliability and trustworthiness.
Acknowledgments
This project was funded by the ERC Starting Grant #101078158 CONVEY and the Israel Science Foundation (ISF) under Grant No. 3079/24.
Appendix A HIGHLIGHTS
“Highlights” Policy Summaries
Our user study uses “Highlights” policy summaries [3] as a baseline. For completeness, we describe the algorithm here. The HIGHLIGHTS algorithm generates an online summary of an agent’s behavior from simulations, using state importance to decide which states to include. A state is considered important if taking a wrong action there significantly decreases future rewards, as determined by the agent’s Q-values. Formally, state importance $I(s)$ is defined as:
$I(s)=\max\limits_{a}Q^{\pi}_{(s,a)}-\min\limits_{a}Q^{\pi}_{(s,a)}.$
HIGHLIGHTS captures trajectories with the most important states encountered in simulations. At each step, it evaluates state importance and adds the state to the summary if its importance exceeds the current minimum in the summary, replacing the least important state. For each state added, it also includes a trajectory of neighboring states and actions.
To address redundancy in similar important scenarios, the HIGHLIGHTS-DIV algorithm extends HIGHLIGHTS by incorporating diversity. HIGHLIGHTS-DIV evaluates a state $s$ by identifying the most similar state $s^{\prime}$ in the summary. It compares $I(s)$ to $I(s^{\prime})$ instead of the minimum importance value. If $I(s)$ is greater, the trajectory including $s^{\prime}$ is replaced with the current trajectory. This approach maintains less important but diverse states, enhancing the information conveyed to users.
Appendix B Information Stored in Experience dataset
Information Description
Episode Number The number of the episode from which the data was collected.
Step Number The specific step within the episode.
Captioned Observation The observation converted into natural language.
Captioned Action The action converted into natural language.
Cumulative Reward The total reward accumulated by the agent up to that step.
Table 5: Description of the data stored in the experience dataset.
Appendix C Summarization Prompt:
[General Instructions]
Generate a focused summary of the RL agent’s policy based on the provided episodes data. Highlight key behaviors, decision-making processes, and patterns specific to this agent. Tailor the summary to reflect unique strategies and actions observed.
Focus on:
• Recurring patterns and behaviors specific to this agent’s policy.
• Detailed analysis of decision-making processes and responses to different stimuli.
• Efficiency in identifying and interacting with relevant objects (e.g., keys, doors).
• Methods used to solve tasks and handle obstacles.
• Comparison of agent’s performance across different episodes.
• Quantitative metrics (e.g., number of steps, success rates) to evaluate efficiency.
• Analysis of navigation strategies and adaptations to the environment.
• Provide insights on the variability and randomness of the agent’s actions and decisions. Look at the distribution of the agent’s actions during the episodes.
• Describe any notable fluctuations or inconsistencies in behavior across episodes.
• Figure out what is the agent’s field of vision and its influence on its decisions and actions, considering it may not see the entire grid but only a part of it.
[Environment Description]
• Goal: Maximize cumulative reward (close to 1) by opening the door.
• The environment is static, with only the agent moving within it. Walls are fixed and form the boundaries of the grid (there are no further obstacles).
• Possible agent’s actions: move forward, turn left, turn right, pick up, toggle, drop.
• Max episode length: 70 steps (failure to open the door if exceeded).
[Interpretation Instructions]
• If the next action is blank, the episode ended successfully. Use the cumulative reward to analyze performance.
• Follow the agent’s received observations to understand its behavior and decision-making process.
• Performance metrics: Number of steps to open the door, cumulative reward.
• Descriptions are from the perspective of the agent, with the number of tiles indicating the distance.
[Output Instructions]
Provide a concise summary focused on specific behaviors and actions of the agent. Formatted for easy understanding by non-technical individuals. Avoid general or trivial statements; focus on specific insights from the agent’s behavior. Include quantitative metrics and comparisons to provide a clear understanding of performance.
[Formatted Experience Dataset]
Appendix D Unlock Goal-Directed Agent Full Summary
This is an example of a summary produced by the SySLLM algorithm, illustrated here using the Unlock goal-directed agent:
The analysis of the RL agent’s behavior across 70 episodes reveals a consistent pattern in its decision-making and interaction with the environment. The agent’s primary goal is to locate and use a key to open a corresponding colored door within a grid environment. Here are the key observations and insights from the agent’s performance:
1. Navigation and Object Interaction: - The agent effectively identifies and navigates towards keys, demonstrating a capability to adjust its path based on the key’s relative position. Once the key is picked up, the agent shifts its focus to finding and reaching the corresponding door. - The agent uses a combination of ’move forward’, ’turn left’, and ’turn right’ actions to navigate towards objectives, showing proficiency in altering its path dynamically based on the environment’s layout and the objects’ positions.
2. Efficiency Metrics: - The agent’s efficiency varies across episodes, with some episodes achieving the goal in fewer steps and higher cumulative rewards. - The average steps taken to complete an episode when successful are approximately 15-25 steps, with cumulative rewards often nearing the maximum, indicating high efficiency in several scenarios.
3. Decision-Making Patterns: - The agent demonstrates a pattern of turning towards the nearest key or door once identified, suggesting a straightforward heuristic of minimizing distance to the target. This behavior is consistent across episodes, indicating a robust underlying policy for basic navigation and objective fulfillment. - In scenarios where the agent is positioned with immediate obstacles or walls, it shows a tendency to turn and reorient itself towards open paths, reflecting an ability to avoid simple traps and dead ends.
4. Handling of Obstacles: - The agent effectively navigates around walls and adjusts its path to reach its targets. There is no evidence of the agent being stuck or failing to navigate around corners, which suggests good spatial awareness and pathfinding within the grid constraints.
5. Variability and Randomness: - The agent’s actions show some variability, particularly in scenarios where multiple paths can lead to the objective. This variability might indicate a level of exploration or randomness in the policy, allowing the agent to discover and adapt to new optimal paths.
6. Performance Consistency: - Across different episodes, the agent’s performance shows consistency in terms of strategy but varies in execution efficiency, likely due to the differing initial conditions and environmental layouts. The agent consistently attempts to pick up the nearest key and unlock the door, aligning with its primary objective.
7. Field of Vision Impact: - The agent’s decisions are heavily influenced by its immediate field of vision, as it tends to react to objects as they become visible. This suggests a limited perception range, requiring the agent to physically move closer to distant objects to detect and interact with them.
In summary, the RL agent is proficient in navigating grid environments to achieve specified goals of picking keys and opening doors. It demonstrates effective use of basic navigational actions, adapts well to the visible environment, and shows a consistent approach to achieving its objectives, albeit with some variability in efficiency possibly due to explorative behaviors or initial positioning. The agent’s performance could potentially be enhanced by expanding its field of vision or optimizing its pathfinding algorithms to increase consistency in achieving higher rewards in fewer steps.
Appendix E Implementation Details
We employed the PPO algorithm from the stable-baselines3 library for our policy network, which takes as input a $K\times K\times 3$ encoded image and a mission string, the latter being encoded using a one-hot scheme. These inputs are combined into a single 2835-dimensional vector. The network architecture features two hidden layers, each comprising 64 neurons, with ReLU activation functions introducing non-linearity. The output layer, designed to match the 6-dimensional action space of the environment, utilizes a softmax activation function to generate a probability distribution over possible actions. Additionally, we normalized the observations. For the short-sighted agent, the observation grid size is $3\times 3\times 3$, while for the goal-directed agent, it is $11\times 11\times 3$.
Hyperparameter Goal-Directed Short-Sighted Dynamic Obstacles Lava Gap Red Blue Doors Crossing
Total Timesteps $2\times 10^{6}$ $1\times 10^{6}$ $2\times 10^{6}$ $2\times 10^{6}$ $2\times 10^{6}$ $3\times 10^{6}$
Number of Environments 8 8 8 16 8 16
Number of Steps 512 512 2048 1024 512 2048
Batch Size 64 64 256 128 64 256
GAE Lambda (gae_lambda) 0.95 0.95 0.95 0.95 0.95 0.95
Discount Factor (gamma) 0.99 0.99 0.99 0.99 0.99 0.99
Number of Epochs 10 10 30 10 10 20
Entropy Coefficient 0.001 0.001 0.01 0.001 0.001 0.01
Learning Rate $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$
Clip Range 0.2 0.2 0.2 0.2 0.2 0.2
Table 6: Hyper-parameters for the PPO algorithm applied to all six agents.
Appendix F Experts instructions:
General Instructions:
Generate a focused summary of the RL agent’s policy based on the provided episodes data. Highlight key behaviors, decision-making processes, and patterns specific to this agent. Tailor the summary to reflect unique strategies and actions observed.
Focus on:
  • Recurring patterns and behaviors specific to this agent’s policy.
  • Detailed analysis of decision-making processes and responses to different stimuli.
  • Efficiency in identifying and interacting with relevant objects (e.g., keys, doors).
  • Methods used to solve tasks and handle obstacles.
  • Comparison of agent’s performance across different episodes.
  • Quantitative metrics (e.g., number of steps, success rates) to evaluate efficiency.
  • Analysis of navigation strategies and adaptations to the environment.
  • Provide insights on the variability and randomness of the agent’s actions and decisions. Look at the distribution of the agent’s actions during the episodes.
  • Describe any notable fluctuations or inconsistencies in behavior across episodes.
  • Figure out what is the agent’s field of vision and its influence on its decisions and actions, considering it may not see the entire grid but only a part of it.
Environment Description:
  • Goal: Maximize cumulative reward (close to 1) by opening the door.
  • The environment is static, with only the agent moving within it. Walls are fixed and form the boundaries of the grid (there are no further obstacles).
  • Possible agent’s actions: move forward, turn left, turn right, pick up, toggle, drop.
  • Max episode length: 70 steps (failure to open the door if exceeded).
Summary Instructions:
The agent description should be at least 100 words. Provide approximately 5 key insights.
Appendix G Scale Used in Task 1
In Task 1 of our study, we utilized a 7-point Likert scale to evaluate participants’ perceptions and understanding of the agent’s behavior as presented in both the video summaries and the natural language summaries. Participants rated their agreement with the following statements, where 1 indicates “Strongly disagree” and 7 indicates “Strongly agree”. The questions were phrased according to the condition—either video or natural language summary.
  1. 1.
    Clarity: “The [video/natural language] summary clearly explained the agent’s actions and decisions shown in the demonstration video.”
  2. 2.
    Understandable: “From the [video/natural language] summary, I understand how the agent’s actions and decisions shown in the demonstration video.”
  3. 3.
    Completeness: “The [video/natural language] summary seemed complete in covering all aspects of the agent’s actions and decisions in the demonstration video.”
  4. 4.
    Satisfaction: “The [video/natural language] summary is satisfying in capturing the agent’s behavior and decisions displayed in the demonstration video.”
  5. 5.
    Useful: “The [video/natural language] summary is useful to my understanding of the agent’s behavior and decisions displayed in the demonstration video.”
  6. 6.
    Accuracy: “The information in the [video/natural language] summary accurately reflected the agent’s behavior and decisions displayed in the demonstration video.”
  7. 7.
    Improvement: “The [video/natural language] summary provides additional insights about the agent’s behavior that are not immediately apparent from watching the demonstration video alone.”
  8. 8.
    Preference: “I prefer receiving information about agent behavior through the [video/natural language] summary rather than just watching the demonstration video.”
These ratings provided quantitative data to assess the effectiveness and clarity of both the video and natural language summaries in conveying the agent’s behavior and decision-making processes. This scale aimed to capture various dimensions of participant satisfaction and understanding, contributing to the overall evaluation of the summaries’ utility in the context of our research.
Appendix H Systematic Exploration of the Prompt Design
The creation of the final prompt was achieved through a structured and iterative exploration process. This process involved a quantitative evaluation of prompt designs based on observed outputs, guided by principles from prompt engineering literature, and tailored to domain-specific requirements. Additionally, the final design was inspired by the Chain of Thought (CoT) [40] prompting paradigm, which encourages models to generate structured, step-by-step reasoning. Below is a detailed breakdown of the methodology used:
Define the Objective
Goal: The primary objective of the prompt was to generate a focused and comprehensive global summary of the policy of the RL agent. The summary needed to highlight key behaviors, decision-making processes, and performance metrics in a manner understandable to both technical and non-technical audiences, while ensuring it could function as a zero-shot prompt without requiring additional training examples.
Key Constraints:
  • The prompt must guide the model to produce specific, concise, and informative summaries.
  • It should minimize general or trivial statements and focus on insights from the agent’s behavior.
Decomposition of Requirements
To meet the objective, the task was broken down into several core components:
  • Behavioral Analysis: Capturing recurring patterns, strategies, and responses to stimuli.
  • Performance Metrics: Including quantitative insights such as success rates and steps taken.
  • Environmental Factors: Reflecting the influence of the agent’s field of vision and static surroundings.
  • Comparison Across Episodes: Addressing variability and randomness in actions.
  • Accessibility: Ensuring the output is clear and digestible for non-technical readers.
Iterative Prompt Design
Initial Prototype:
  • Focused on general instructions for summarization.
  • Included high-level tasks such as “describe the agent’s behavior” without specifying details.
Issues Identified:
  • Outputs were overly generic, lacked depth, and failed to focus on specific behaviors or metrics.
Refinement 1: Add Specific Focus Areas
  • Incorporated bullet points to guide the model to focus on particular aspects, such as “recurring patterns,” “quantitative metrics,” and “navigation strategies.”
Observations:
  • Improved relevance and depth of the summaries.
  • However, the outputs lacked consistency in formatting and interpretability.
Refinement 2: Structured Prompt Sections
  • Segmented the prompt into distinct parts:
    • General Instructions
    • Environment Description
    • Interpretation Instructions
    • Output Instructions
    • Formatted Experience Dataset
Observations:
  • Enhanced structure improved consistency.
  • More detailed context in “Environment Description” provided clarity for the model to ground its responses.
Refinement 3: Inspired by Chain of Thought (CoT) Reasoning
  • The prompt was designed to encourage a step-by-step analysis, mirroring the CoT paradigm:
    • Each bullet point and section was treated as a sub-task requiring focused attention.
    • For example, instructions like “Analyze navigation strategies and adaptations to the environment” explicitly directed the model to break down its reasoning into smaller, manageable steps.
Observations:
  • Outputs exhibited improved logical flow and comprehensive coverage of required aspects.
  • The structured approach mitigated issues with overly generic or shallow responses.
Refinement 4: Emphasize Quantitative and Comparative Analysis
  • Added explicit instructions to include metrics like “number of steps” and “success rates.”
  • Introduced the requirement to compare the agent’s performance across episodes.
Observations:
  • Summaries became more data-driven and analytical.
  • Increased attention to variations in the agent’s behavior.
Refinement 5: Addressing Accessibility
  • Adjusted language in the “Output Instructions” to ensure summaries were understandable to non-technical audiences.
  • Included a directive to avoid trivial statements.
Final Testing:
  • Conducted multiple test runs with varied episode datasets.
  • Evaluated the prompt’s ability to guide the model toward producing outputs that met the objective.
  • Fine-tuned phrasing for clarity and focus.
Key Design Considerations
Clarity and Specificity:
  • Each section of the prompt was crafted to minimize ambiguity, ensuring the model understood the task requirements.
Structure Inspired by CoT:
  • The step-by-step breakdown mirrored the CoT prompting approach, which is known to improve reasoning and response quality in large language models.
Focus on Insightful Analysis:
  • By explicitly asking for “variability,” “distribution of actions,” and “quantitative comparisons,” the prompt steered the model toward generating meaningful insights.
Evaluation and Lessons Learned
Evaluation:
  • Outputs were analyzed for relevance, specificity, and clarity.
  • Feedback from test runs informed iterative improvements.
Lessons Learned:
  • Prompts benefit from structured sections that provide clear and detailed guidance.
  • Incorporating CoT-inspired design principles encourages logical, step-by-step reasoning in outputs.
  • Tailoring language for accessibility improves utility for non-technical audiences.
Rationale for the Final Design
The final prompt integrates the following elements:
  • Comprehensive Instructions: Ensuring detailed and targeted outputs.
  • Quantitative Focus: Providing measurable insights for evaluating agent performance.
  • Clarity and Accessibility: Catering to a broad audience, including non-technical users.
  • Structure Inspired by CoT: Encouraging the model to follow a logical sequence in generating summaries.
This systematic process, incorporating insights from the Chain of Thought paradigm, demonstrates the thoughtful process taken to ensure the prompt is both effective and robust for summarizing RL agent policies.

### v2

@@
-SySLLM: Generating Synthesized Policy Summaries for Reinforcement Learning Agents Using Large Language Models
+From Actions to Words: Towards Abstractive-Textual Policy Summarization in RL
 Abstract
-Policies generated by Reinforcement Learning (RL) algorithms can be difficult to describe to users, as they result from the interplay between complex reward structures and neural network-based representations.
-This combination often leads to unpredictable behaviors, making policies challenging to analyze and posing significant obstacles to fostering human trust in real-world applications.
-Global policy summarization methods aim to describe agent behavior through a demonstration of actions in a subset of world-states.
-However, users can only watch a limited number of demonstrations, restricting their understanding of policies.
-Moreover, those methods overly rely on user interpretation, as they do not synthesize observations into coherent patterns.
-In this work, we present SySLLM (Synthesized Summary using LLMs), a novel method that employs synthesis summarization, utilizing large language models’ (LLMs) extensive world knowledge and ability to capture patterns, to generate textual summaries of policies.
-Specifically, an expert evaluation demonstrates that the proposed approach generates summaries that capture the main insights generated by experts while not resulting in significant hallucinations.
-Additionally, a user study shows that SySLLM summaries are preferred over demonstration-based policy summaries and match or surpass their performance in objective agent identification tasks.
-[^1] All relevant resources for this study, such as SySLLM summaries, HIGHLIGHTS summaries, expert summaries, experience videos, the source code, and other materials are available at https://osf.io/c78jb/?view_only=76fbf59131d4409bac222e5929a654f2.
+Policies generated by Reinforcement Learning (RL) algorithms are difficult to explain to users, as they emerge from the interaction of complex reward structures and neural network representations.
+Consequently, analyzing and predicting agent behavior can be challenging, undermining user trust in real-world applications.
+To facilitate user understanding, current methods for global policy summarization typically rely on videos that demonstrate agent behavior in a subset of world states.
+However, users can only watch a limited number of demonstrations, constraining their understanding.
+Moreover, these methods place the burden of interpretation on users by presenting raw behaviors rather than synthesizing them into coherent patterns.
+To resolve these issues, we introduce SySLLM (Synthesized Summary using Large Language Models), advocating for a new paradigm of abstractive-textual policy explanations.
+By leveraging Large Language Models (LLMs)—which possess extensive world knowledge and pattern synthesis capabilities—SySLLM generates textual summaries that provide structured and comprehensible explanations of agent policies.
+SySLLM demonstrates that LLMs can interpret spatio-temporally structured descriptions of state-action trajectories from an RL agent and generate valuable policy insights in a zero-shot setting, without any prior knowledge or fine-tuning.
+Our evaluation shows that SySLLM captures key insights, such as goal preferences and exploration strategies, that were also identified by human experts.
+Furthermore, in a large-scale user study (with 200 participants), SySLLM summaries were preferred over demonstration-based summaries (HIGHLIGHTS) by a clear majority (75.5%) of participants.
+journal: Artificial Intelligence
+[…]
+[…]
 1 Introduction
-Explainable Reinforcement Learning (XRL) methods aim to enable human users to understand agents’ policies and decision-making processes.
-Such understanding is crucial for fostering trust and ensuring the real-world deployment of RL agents.
-Several approaches to XRL have been developed in recent years [25], including saliency maps that visualize the agent’s attention, reward bars that break down expected utility into components, and causal explanations that clarify why actions were taken.
-Other methods extract logical rules to describe when actions are taken, but these approaches rely on predefined logical representations and often focus on individual actions rather than broader behavior patterns.
-Despite these advancements, experiments reveal that human users still struggle to form accurate mental models of agent behavior [6, 19, 23].
-Recent advancements in large, pre-trained language models (LLMs) present a novel opportunity to enhance explanation methods for reinforcement learning (RL) policies.
-These models exhibit exceptional in-context learning abilities and possess extensive commonsense knowledge about human behavior, which makes them promising tools for generating meaningful, language-based explanations of RL policies.
-LLMs have already been employed in constructing world models and planning with domain-independent planners, showcasing their potential for reasoning and providing explanation [14].
+Reinforcement learning (RL) agents are increasingly deployed in high-stakes domains, yet their opaque decision-making processes remain a barrier to trust and adoption.
+Explainable RL (XRL) methods aim to bridge this gap by surfacing interpretable accounts of agent behavior.
+Despite this promise, current XRL methods face significant limitations.
+Saliency maps [1, 2, 3, 4] highlight critical states but offload pattern discovery to users.
+Trajectory demonstrations [5, 6, 7, 8] constrain the understanding to cherry-picked scenarios.
+Rule extraction methods [9, 10, 11, 12], while providing structured representations, often struggle to capture the full complexity of learned policies.
+Many of these approaches share a common challenge: they present transformed observations rather than synthesized insights, requiring users to manually extrapolate agent intent, adaptability, and failure modes from fragmented evidence [13].
+Recent advances in large language models (LLMs) offer a transformative alternative.
+With their capacity for abstraction, commonsense reasoning, and natural language generation [14], LLMs hold unique potential to synthesize agent trajectories into human-interpretable narratives.
+However, this promise comes with a challenge, LLMs’ static text-based training contrasts with the dynamic, spatiotemporal nature of reinforcement learning environments [15].
+Early applications of LLMs in RL, such as world modeling [16], hint at their promise, but generating faithful global policy summaries—cohesive accounts that generalize beyond isolated states—remains unexplored.
 Figure 1: Collecting the experience dataset (Section 4.3).
-In this paper, we aim to leverage LLMs to generate comprehensive global explanations of RL agent policies.
-Specifically, our goal is to use LLMs to produce natural language summaries that encapsulate key patterns of behavior, capabilities, and limitations of RL policies.
-This approach seeks to build upon previous work that utilized demonstration-based policy summaries [4], where agent policies are described through selected trajectories in video.
-However, these methods are inherently constrained by the representativeness of the selected scenarios and the limited number of demonstrations that can feasibly be observed.
-They do not synthesize observations into coherent summaries, leaving users to rely heavily on their own interpretations to discern patterns.
-In contrast, we hypothesize that LLM-generated summaries can effectively distill these patterns into concise, coherent narratives, providing users with a clearer and more accessible understanding of agent policies.
-Leveraging LLMs to generate accurate and comprehensible policy summaries is a non-trivial task.
-While LLMs exhibit remarkable reasoning abilities, they are prone to hallucinations—producing outputs that seem plausible but deviate from the agent’s true behavior [16].
-This raises questions about whether such descriptions can reliably reflect the nuances of RL agent policies.
-Moreover, their utility heavily depends on end users’ ability to interpret them.
-Laypeople, who often lack familiarity with RL concepts, may struggle to form accurate mental models of agent behavior.
-Natural language explanations, while accessible, risk oversimplifying or failing to convey subtle patterns [6].
-A further challenge lies in capturing the spatial and temporal relationships inherent in RL environments.
-These relationships evolve continuously over many steps and episodes: spatial dependencies include changing object positions, while temporal dependencies arise from sequential actions.
-Since LLMs are trained primarily on static text, they may struggle to model these dynamics effectively [24].
-Together, these factors complicate the task of distilling and communicating agent behaviors in a way that remains faithful to the policy while being accessible to users.
-Our method, Synthesized Summary using LLMs (SySLLM), leverages LLMs to generate comprehensive summaries of RL agent behavior.
-The proposed pipeline first collects the agent’s interactions within the environment (Figure 1) and converts these into natural language descriptions of observations and actions.
-These descriptions are then used to prompt an LLM, which synthesizes a natural language summary that captures key patterns in the agent’s behavior (Figure 2).
-We implement this methodology across five distinct environments, generating language-based summaries for agents with diverse behavioral patterns.
-To evaluate SySLLM, we employ two complementary methods.
-First, we compare its summaries to those created by six RL experts, finding substantial agreement between the two.
-Second, we conduct a user study where participants interact with two types of policy summaries: visual summaries generated by the HIGHLIGHTS-DIV algorithm [3] and language-based summaries generated by SySLLM.
-Participants rated SySLLM higher in subjective preferences and performed equally well or better at identifying agents based on these summaries.
-These results highlight the effectiveness of SySLLM in both comprehensibility and practical utility.
-The main contributions of our work are:
-  • We introduce Synthesized Summary using LLMs (SySLLM), a novel and versatile method that generates holistic, synthesized, language-based summaries of RL agent policies using LLMs.
-  • We validate the correctness of SySLLM by comparing its summaries to those of RL experts, showing substantial alignment and demonstrating that SySLLM accurately captures key policy details.
-  • We conduct a large-scale user study, demonstrating that participants prefer SySLLM over a state-of-the-art demonstration-based policy summarization method in a summary preference task and perform equally well or better in an agent identification task.
+In this work, we advocate for an abstractive-textual approach for policy summarization.
+We introduce SySLLM, a framework that leverages LLMs to generate holistic summaries of RL policies by systematically distilling agent trajectories into structured narratives.
+SySLLM addresses the limitations of previous methods through a two-stage process: (1) converting raw agent-environment interactions into temporally grounded language descriptions (Fig.
+1), and (2) prompting LLMs to synthesize these into coherent summaries that capture behavioral patterns, strategic tendencies, and environment-specific adaptations (Fig.
+2).
+SySLLM performs impressively well—it captures agent preferences (as induced by different reward functions), identifies sub-optimal behavior in under-trained agents, and distinguishes between diverse exploration strategies.
+Beyond summarizing the agent’s behavior, SySLLM offers insights into the effectiveness of the agent’s strategy.
+SySLLM demonstrates that LLMs can interpret spatiotemporally structured descriptions of state-action trajectories from an RL agent.
+While these descriptions are written in natural language, they differ significantly from the kinds of text LLMs are typically trained on, such as stories or dialogue.
+Instead, they represent a distinct modality: sequential, decision-driven behavior grounded in the dynamics of an environment.
+Despite this, SySLLM shows that LLMs can process such unfamiliar input and, in a zero-shot setting—without any prior knowledge or fine-tuning—generate valuable policy insights.
+Our framework is validated through multi-faceted evaluation.
+First, comparisons with RL experts reveal that SySLLM summaries align closely with expert annotations, capturing nuanced behaviors that demonstration-based methods do not explicitly highlight.
+Second, a large-scale user study shows that participants prefer SySLLM over state-of-the-art visual summaries such as HIGHLIGHTS [17], while also achieving equal or better accuracy in agent identification tasks.
+In summary, our main contributions are the following:
+  • We introduce SySLLM, a framework that takes a step toward generating abstractive-textual explanations of RL policies using LLMs.
+SySLLM synthesizes structured narratives that reveal agent intent, adaptability, and limitations by systematically analyzing agent trajectories, enabling users to grasp these behaviors without manual pattern matching.
+SySLLM eases the burden of interpreting agent behavior by digesting complex dynamics into coherent summaries.
+To the best of our knowledge, this is the first work to produce global, abstractive explanations of an RL agent’s policy in natural language.
+  • Generated summaries align with RL experts.
+We evaluate the effectiveness of SySLLM by comparing its summaries with those written by RL experts, demonstrating substantial alignment and accurate coverage of key policy details.
+To evaluate faithfulness—an area often prone to issues in automatic summarization [18, 19]—we additionally compute precision scores.
+Our results show that SySLLM maintains high precision overall, indicating that its summaries rarely introduce unsupported claims.
+  • Users prefer our textual summaries.
+We conduct a large-scale user study with 200 participants showing that participants clearly prefer SySLLM over a leading demonstration-based policy summarization method on a summary preference task (75.5%), while performing equally well or better in an agent identification task.
 Figure 2: Generating natural language global policy summaries (Section 4.4).
 2 Background and Related Work
-Explainable RL (XRL)
-Our work extends the body of research concerned with explaining RL policies and particularly the line of work on global policy explanations (see Amitai and Amir [5] for a recent survey).
-Policy summarization approaches present users with demonstrations of the agent acting in the environment to convey its behavior.
-Several criteria for extracting such summaries have been proposed in the literature, including state importance, diversity, visitation frequency, and policy reconstruction [3, 17, 23, 18, 34].
-While these approaches were shown to improve people’s ability to assess agent capabilities and predict agent behavior, users can only grasp a limited amount of demonstrations, and experiments show that their interpretation of policies highly depends on the subset of examples shown to them [23].
-Other works propose learning a simplified model of the policy (e.g., a decision tree or a set of logical rules) that captures the essence of the agent’s decision-making [38, 8, 7].
-Yet, these approaches are also limited due to the tradeoff between the accuracy of the simplified model and its interpretability to users.
-We hypothesize that LLMs may offer a complementary approach to global explanations by capturing meaningful patterns of agent behavior and presenting them in understandable language to users.
-Pre-trained Language Models in RL
-Large Language Models (LLMs) are probabilistic text models trained on substantial text datasets.
-Models like GPT [30], which are autoregressive, excel in text generation by sequentially predicting future words based on the preceding text.
-Such models can be useful in the context of RL as they encode common-sense knowledge about the world.
-Conversely, models such as BERT [12], which are based on the masked framework, focus on understanding deep sentence semantics by training to recognize parts of text that are intentionally obscured.
-Encoder-decoder models merge these traits to adeptly handle tasks requiring both the interpretation and creation of text.
-These models have been extensively pre-trained on large-scale text corpora, equipping them with the ability to tackle complex language processing tasks with minimal specific instruction, often achieving impressive results [9].
-This extensive training allows them to efficiently process and summarize complex narratives, and to identify relevant patterns within large datasets.
-Recent work has developed methods that use LLMs in the training of RL policies.
-For example, LLMs are being used to guide intrinsic exploration by suggesting plausibly useful goals and providing intrinsic rewards [13, 27].
-Other works use LLMs to generate useful state abstractions [29, 28].
-Additionally, LLMs are employed to provide natural language instructions and dynamic feedback, enhancing agent learning in complex, open-ended environments [39].
-Differently from these approaches, in this work, we aim to use LLMs to explain RL policies.
-3 Problem Statement
+2.1 Explainable Reinforcement Learning
+Explainable reinforcement learning (XRL) aims to make the behavior of RL agents more understandable to human users.
+Traditional XRL techniques fall into several major categories: saliency-based methods, demonstration-based summaries, surrogate models, and causal explanations [20].
+Saliency and Visualization Techniques
+Early XRL work often adopted visualization tools such as saliency maps [1, 2, 3, 4] to indicate which parts of the input space influenced the agent’s decisions.
+While these tools provide some insight, they require users to manually interpret the spatial and temporal relevance of highlighted regions, leading to potential misinterpretation.
+Studies show these methods are local and fragile [21], providing little insight into the long-term policy structure.
+Demonstration-Based Summaries
+Another common approach involves selecting representative episodes or critical states that demonstrate agent behavior [5, 6, 7, 8].
+These methods allow users to observe behavior directly, but rely heavily on cherry-picked scenarios, with high cognitive load placed on the user to infer strategic intent from limited data.
+Work by [22] shows that users often struggle to generalize from such demonstrations, particularly when the agent behavior is stochastic.
+Surrogate Models
+An alternative strategy uses model distillation [23] to learn interpretable approximations of agent behavior, such as decision trees [24, 25].
+These methods strike a balance between interpretability and fidelity: overly simple models fail to capture nuances of the agent’s strategy, while complex surrogates become as opaque as the original policy.
+Recent works explore bounding MDPs [26] or abstract policy graphs [27] to improve this trade-off, but scalability remains an issue.
+Causal Explanations
+A complementary direction models the agent’s reasoning through structural causal models or influence graphs [28, 29].
+These allow users to query “why” an agent took a specific action and offer contrastive explanations that align with how humans seek understanding.
+However, such models often require significant prior knowledge of the domain or agent internals, limiting their applicability.
+Reward Decomposition
+Other methods explain agent behavior by breaking down the overall reward into meaningful components, such as safety, efficiency, or task-specific objectives.
+This allows explanations that clarify how each reward type influences the agent’s decisions and highlights trade-offs between competing objectives [30, 31, 32].
+These approaches help users understand agent priorities and decision rationale by making the contribution of each reward component explicit.
+2.2 Large Language Models in Reinforcement Learning
+Pre-trained Large Language Models (LLMs) have revolutionized NLP through their ability to distill patterns from vast datasets into coherent text.
+Autoregressive models like GPT [33] generate human-like narratives by predicting token sequences, while encoder-decoder architectures excel at summarizing complex inputs [34].
+Those recent advances have introduced new possibilities for interacting with and explaining RL agents.
+While much prior work integrates LLMs into the training process e.g., guiding exploration [35] or building world models [16], less attention has been paid to their potential as explanation generators.
+[36] explore using LLMs to generate local post-hoc explanations of individual decisions based on distilled behavior representations, but such approaches do not aim to provide global summaries of agent behavior.
+Trajectory Annotation and Feedback
+Some works use LLMs to narrate agent behavior in real-time or provide textual feedback.
+[37, 38] demonstrate that LLMs can be prompted to interpret actions or suggest improvements mid-episode.
+These studies show promise for LLM-based commentary, though they often lack grounding in actual agent dynamics, leading to inconsistencies.
+Model-Based Reasoning and Simulation
+In world modeling applications, LLMs have been used to generate or refine symbolic models of environment dynamics [39].
+Here, the LLM acts as a high-level planner or simulator, encoding causal knowledge that guides behavior.
+This further validates the hypothesis that LLMs can represent abstract behavioral structures, which is essential for high-level summarization.
+These applications largely treat LLMs as agents or environment models, not as explainers.
+When LLMs are used for explanation, they must be carefully grounded in observed experience to avoid plausible-sounding but inaccurate generalizations [40].
+2.3 Natural Language Explanations for RL Agents
+The use of natural language as a modality for RL explanations offers distinct advantages over visual or rule-based approaches.
+Language allows for abstraction, temporal coherence, and commonsense reasoning—all of which are challenging to express in saliency maps or decision trees.
+Template-Based Explanations
+Early works, such as [41], demonstrated that combining learned MDP structures with templated natural language generation could enable contrastive explanations (e.g., “Why didn’t you do X?”) through predicate-based state abstraction, allowing generalization across policy architectures.
+Later, simpler approaches reemerged, such as [10], which employed rigid, static templates to directly translate policies into interpretable statements (e.g., “If near wall, turn left”), prioritizing accessibility but inheriting brittleness and limited expressivity.
+More recent systems like [42] also rely on predefined action-condition mappings, facing similar limitations.
+Rationalization and Free-Form Generation
+Subsequent work adopted neural methods to generate textual rationales.
+[43] introduced AI rationalization training sequence-to-sequence models to translate trajectories into natural language.
+Later, [44] introduced graph-based policy summaries that enhance user understanding by clustering states and abstracting the state space.
+While those natural language explanations methods offer interpretability, they are inherently limited by their reliance on handcrafted rules and rigid structures, making them brittle and unable to express nuanced or unexpected behaviors.
+Rationalization methods, which translate individual trajectories into free-form text, tend to focus on local justifications rather than providing a holistic view of agent behavior.
+These approaches often lack scalability across diverse scenarios and struggle to generalize patterns observed over many episodes.
+Moreover, they typically do not leverage broader world knowledge or incorporate high-level reasoning, which can limit the clarity, depth, and utility of the resulting explanations for end users.
+3 Motivation for Abstractive-Textual Explanations
+Despite recent advances in explainable reinforcement learning (XRL), existing methods face a persistent trade-off between expressiveness, scalability, and accessibility.
+Extractive approaches—such as saliency maps or curated demonstrations—highlight fragments of agent behavior but often fail to integrate them into coherent, high-level narratives.
+Symbolic models, including logical rules and decision trees, offer structural clarity but require handcrafted features and often struggle to scale in complex or high-dimensional settings.
+These limitations point to a need for explanation methods that can synthesize global behavioral insights—capturing patterns, adaptation, and strategy over time—without relying on structured input representations or human-crafted abstractions.
+We advocate for a textual-abstractive paradigm in XRL, where natural language is used to summarize agent behavior across episodes.
+Text offers a flexible medium for conveying temporal dependencies, recurring behaviors, and decision-making logic in a format accessible to non-experts.
+Unlike extractive or symbolic methods, textual explanations can operate over raw behavioral data without predefined rules or model internals, making them broadly applicable across domains.
+Recent advances in large language models (LLMs) open new opportunities for explainability in reinforcement learning.
+With SySLLM, we introduce a textual-abstractive paradigm that uses natural language to summarize agent behavior over time—capturing patterns, strategic intent, and adaptation.
+In this work, we show that LLMs are particularly well-suited for this approach, when guided by an appropriate framework, they can recognize sequential structure and express it in fluent, context-aware language.
+We view this paradigm as a holistic and scalable complement to existing XRL methods, and as part of a broader shift toward language-based interfaces for interpreting learned behavior.
+4 SySLLM Method
+Our SySLLM (Synthesized Summary using LLMs) method consists of several key components, which we formalize in Algorithm 1.
+In the first phase (illustrated in Figure 1), we run the agent in its environment and perform a captioning procedure to convert experiences into natural language.
+These captioned experiences are logged into an experience dataset (Sections 4.2 and 4.3).
+In the second phase, which we illustrate in Figure 2, we take the experience dataset as input and output a natural language summary that aims to capture patterns in the agent’s policy (Section 4.4).
+To do so, we compose a prompt with the formatted experience dataset that is then sent to an LLM to generate the summary.
+4.1 Problem Statement
 We define a partially observed Markov Decision Process (POMDP) as a tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{O},T,O,R,\gamma\rangle$, where observations $o\in\mathcal{O}$ are derived from states $s\in\mathcal{S}$ and actions $a\in\mathcal{A}$ via $O(o|s,a)$.
@@
 Formally, a summarization is a mapping from $(\mathcal{M},\Pi)$ to $\mathcal{T}$, where $(\mathcal{M},\Pi)$ represents environment-policy pairs, and $\mathcal{T}$ is the set of natural language summaries.
-4 Methodology
-Our method consists of several key components, which we formalize in Algorithm 1.
-In the first phase (illustrated in Figure 1), we run the agent in its environment and perform a captioning procedure to convert experiences into natural language.
-These captioned experiences are logged into an experience dataset (Sections 4.1 and 4.2).
-In the second phase, which we illustrate in Figure 2, we take the experience dataset as input and output a natural language summary that aims to capture patterns in the agent’s policy (Section 4.3).
-To do so, we compose a prompt with the formatted experience dataset that is then sent to an LLM to generate the summary.
 4.2 Captioner
@@
 The for loop (Line 2) iterates over multiple episodes, initializing the step counter $t=1$ at the beginning of each episode (Line 3) and obtaining the initial observation $o_{1}$ (Line 4).
-The variables $epReward$ and $done$ are also initialized to track the cumulative reward and episode completion status, respectively (Lines 56).
+The variables $epReward$ and donedone are also initialized to track the cumulative reward and episode completion status, respectively (Lines 56).
 The while loop (Line 7) continues until the episode concludes.
@@
 Subsequently, the environment is updated using the selected action, and the next observation is retrieved (Line 10).
-Finally, the cumulative reward and step index are updated accordingly (Lines 1112). The information stored in the experience dataset is detailed in Section B of the Appendix.
+Finally, the cumulative reward and step index are updated accordingly (Lines 1112).
+The information stored in the experience dataset is detailed in B.
 4.4 Generating Natural Language Summaries
@@
 This format captures the sequential and contextual details of the agent’s interactions, creating a comprehensive narrative of its decision-making process (Line 15).
-The prompt design, inspired by Chain-of-Thought reasoning [40], decomposes complex behavioral analysis into specific components.
+The prompt design, inspired by Chain-of-Thought reasoning [47, 48], decomposes complex behavioral analysis into specific components.
 By incorporating detailed environmental parameters and quantitative evaluation guidelines, the prompt effectively guides the model’s analysis, enhancing its ability to generate insightful summaries of reinforcement learning behavior.
-This design was refined through iterative exploration, as detailed in Section H of the Appendix.
+This design was refined through iterative exploration, as detailed in I.
 The prompt structure consists of four main components:
@@
 The LLM then produces the summary based on the prompt and the formatted experience dataset (Line 16).
-Algorithm 1 SySLLM algorithm
+Algorithm 1 SySLLM
 Input: Environment $ENV$, trained policy $\pi$, captioners $C_{\text{observation}}$, $C_{\text{action}}$
@@
 Output: summary $\in\mathcal{T}$
-1:  Initialize experience dataset $ED$
-2:  for $i=1$ to $N$ do
+1: Initialize experience dataset EDED
+2:for $i=1$ to NN do
 3:  Let $t\leftarrow 1$
-4:     $o_{t}\leftarrow$ $ENV$.RESET()
+4:  $o_{t}\leftarrow$ ENVENV.RESET()
 5:  $epReward\leftarrow 0$
@@
 8:   $a_{t}\sim\pi(\cdot\mid o_{t})$
-9:        $ED$.ADD($C_{\text{observation}}(o_{t})$, $C_{\text{action}}(a_{t})$, $epReward$)
-10:        $o_{t+1},r_{t+1},done\leftarrow$ $ENV$.STEP($a_{t}$)
+[…]
+10:   $o_{t+1},r_{t+1},done\leftarrow$ ENVENV.STEP($a_{t}$)
 11:   $epReward\leftarrow epReward+r_{t+1}$
@@
 14:end for
-15:  $formattedInput$leftarrow $P$ + FORMAT($ED$)
-16:  $summary$leftarrow LLM($formattedInput$)
-17:  return $summary$
+15:formattedInputformattedInput $\leftarrow$ $P$ + FORMAT(EDED)
+16:summarysummary $\leftarrow$ LLM(formattedInputformattedInput)
+17:return summarysummary
 5 Implementation
-We demonstrate our approach in five diverse environments to evaluate summaries for seven agents, each with a unique policy. Specifically, we use the MiniGrid framework [10].
-5.1 Captioners
+We showcase our method in five distinct environments from the MiniGrid Framework [49] as well as the challenging Crafter environment [50].
+5.1 MiniGrid
+MiniGrid is a framework for 2D based environments that feature goal-oriented tasks where agents navigate mazes, interact with objects like doors or keys, and solve challenges.
+We evaluate our approach in five distinct MiniGrid environments using seven different agents, each following a unique policy.
+5.1.1 Captioners
 To convert raw observations and actions into natural language, we designed a rule-based captioning system for the agent’s grid-based representation of the environment.
@@
 We evaluated the effectiveness of this captioner by experimenting with and refining its design, ensuring the generated descriptions are accurate and comprehensive.
-5.2 Agents’ Policy Training
+5.1.2 Agents’ Policy Training
 We trained three qualitatively different agents for the MiniGrid-Unlock environment:
@@
 Using SySLLM, we generated a policy summary for each of these agents.
-All of the agents are trained using the PPO algorithm [33, 31].
-An overview of agent performance is shown in Table 1.
+All of the MiniGrid agents are trained using the PPO algorithm [51, 52].
+An overview of agents’ performance is shown in Table 1.
 The mean reward for the first three agents in the Unlock environment shows a declining trend, which aligns with the policies that each of these agents employ.
 In addition, the success rate of the agents in the other environments remains consistently high.
-Implementation details are provided in Section E of the Appendix.
+Implementation details are provided in F.
 Agent Mean Reward ± SD Mean Length Success Rate
@@
 Dynamic Obstacles 0.78 ± 0.06 17.20 1.00
-Lava Gap 0.896 ± 0.02 10.82 1.00
+Lava Gap 0.90 ± 0.02 10.82 1.00
 Red Blue Doors 0.70 ± 0.26 17.06 0.88
 Crossing 0.67 ± 0.18 24.80 0.94
-Table 1: Report of agents’ performance metrics for 500 episodes.
-5.3 Generating Policy Summaries
+Table 1: Report of agents’ performance metrics for 500 episodes for the MiniGrid environments across 3 different seeds.
+5.1.3 Generating Policy Summaries
 For each agent, we collect an experience dataset that logs the captioned observations, actions taken by the agent, and additional relevant data.
@@
 We employed the gpt-4-turbo model with a temperature of $0.2$ [53].
-The prompt used is shown in Section C of the Appendix.
+The prompt used is shown in C.
 In selecting the final summary, we employed a methodical approach by first generating ten distinct summaries through Line 16 of our algorithm.
@@
 Summaries closer to the centroid tend to include more general statements that apply broadly across different scenarios, while those further away incorporate more specific details unique to particular contexts.
-An example of a selected summary is provided in Section D of the Appendix.
+An example of a selected summary is provided in D.
+5.2 Crafter
+Secondly, we evaluate SySLLM in the Crafter environment, a 2D adaptation of Minecraft.
+Similar to Minecraft, Crafter features a procedurally generated, partially observable world where agents can gather resources and craft artifacts structured within an achievement tree.
+This tree outlines all possible achievements along with their prerequisites.
+Crafter lacks a singular main objective, allowing agent progress to be measured by the achievements it unlocks.
+5.2.1 Captioners
+To convert the environment observation representations and actions in Crafter into natural-language format, we implemented a captioning system tailored to the game’s mechanics.
+The Observation Captioner component generates comprehensive descriptions of the player’s current state, including inventory contents, achievements, and spatial relationships to nearby resources and threats.
+For example, a typical observation might be transformed from raw data into ”The player is facing right.
+Inventory: 2 wood, 1 stone.
+A zombie is ahead…”.
+The spatial awareness system maps objects to their relative coordinates.
+The Action Captioner component complements this by translating action indices into meaningful descriptions like “place_table”, “move_right” or “make_wood_sword”, providing context for the agent’s decisions.
+Figure 3 provides an example of how the captioners translate the agent’s trajectory into natural language descriptions.
+Figure 3: This figure illustrates four steps from a trajectory of the Resource-Collector agent in the Crafter environment, alongside their corresponding natural language captions generated using the observation and action captioners, $C_{\text{observation}}$ and $C_{\text{action}}$.
+For each step, the captions describes the agent’s inventory status, the object currently in front of it, and the next action selected by the agent.
+A textual representation of the visible grid (highlighted in blue) is also included to reflect the agent’s local perception.
+Additionally, all unique achievements unlocked by the agent throughout the trajectory are summarized in red.
+5.2.2 Agents’ Policy Training
+We trained two qualitatively different agents for the Crafter environment.
+The first, named Resource-collector Agent, is a well-trained agent capable of navigating in Crafter with efficiency and skill.
+The agent is capable to search for food and water, defend against monsters, collect materials, and craft tools while unlocking semantically meaningful achievements along the achievement tree.
+This agent was trained using the DreamerV3 algorithm [54].
+The second agent, a Random Agent, selects actions uniformly at random from the available action space without any learned strategy.
+Implementation details are provided in F.
+5.2.3 Generating Policy Summaries
+Similar to the MiniGrid environments, we gather an experience dataset for each agent, recording captioned observations, the actions taken by the agent, and other relevant information.
+Specifically, we log five episodes per agent.
+Due to context window limitations, each episode is summarized individually, and the resulting episode summaries are then combined into a single comprehensive summary.
+The prompts used are shown in E.
 5.3 Insights From Summaries
-Using the SySLLM method, we generated summaries capturing key aspects of each agent’s decision-making and behavior.
-This section presents key insights from these summaries, concisely illustrating agents’ policies and critical action patterns.
-Table 2 provides a overview of these insights, showcasing their alignment with task objectives and how agents behave across various environments.
+Using the SySLLM method, we produced summaries that highlight essential aspects of each agent’s decision-making and behavior.
+This section outlines key insights derived from these summaries, offering a clear depiction of the agents’ policies and significant action patterns.
+Table 3 presents an overview of these insights for the MiniGrid agents, while Table 4 provides an overview for the Crafter agents.
+These insights showcase the alignment of SySLLM-generated summaries with task objectives and illustrate how agents perform across diverse environments.
 For the goal-directed agent in the Unlock environment, Insight (2) states that the agent “demonstrates a pattern of turning towards the nearest key or door once identified, suggesting a straightforward heuristic of minimizing distance to the target.” This reflects the agent’s efficient and consistent behavior, dynamically adjusting its path to achieve objectives with minimal steps.
@@
 These observations demonstrate SySLLM’s ability to provide a nuanced understanding of both the agent’s strengths in straightforward scenarios and its struggles in more complex configurations.
+For the Resource-collector agent in the Crafter environment, SySLLM effectively captures the agent’s behavior as a resource-collector, emphasizing its focus on gathering essential materials.
+Insight (1) highlights that “The agent consistently prioritizes acquiring fundamental resources such as wood, stone, and drink, which are critical for crafting tools and sustaining basic survival metrics.” Furthermore, SySLLM captures the agent’s achievement-oriented tendencies while identifying limitations in unlocking advanced achievements.
+As stated in Insight (4), “The agent reliably achieves milestones related to resource collection and basic tool crafting but demonstrates difficulty in attaining more advanced achievements, suggesting potential areas for improvement.” These observations are validated by performance metrics presented in Table 2, which report a mean reward of 10.42—indicating relatively high performance—and a mean of 11.33 unique achievements unlocked, reinforcing the agent’s proficiency in foundational tasks while highlighting room for growth in advanced capabilities.
+Agent Mean Reward ± SD Mean Length Mean Unique Achievements
+Resource-Collector 10.43 ± 2.11 234.6 11.33
+Random 1.39 ± 1.19 164.4 2.29
+Table 2: Report of agents’ performance metrics for 100 episodes for the Crafter environment across 3 different seeds.
 Agent Insights from Summaries
 Unlock
-Goal-Directed
+Goal–
+directed
 (1) The agent effectively identifies keys and adjusts its path based on their relative position, shifting focus to unlocking the door.
 (2) It consistently turns towards the nearest key or door, minimizing distance, which remains consistent across episodes.
-(3) The agent completes episodes efficiently, averaging 15-25 steps with near-maximal cumulative rewards.
+(3) The agent completes episodes efficiently, averaging 15–25 steps with near-maximal cumulative rewards.
 Unlock
-Short-sighted
+Short–
+sighted
 (1) The agent follows a right-wall method, moving forward until encountering an obstacle before turning.
@@
 (3) Decisions are heavily influenced by its immediate field of vision, reacting only to nearby objects.
-Unlock Random (1) The agent exhibits unstructured behavior, often repeating unnecessary actions.
+Unlock
+Random
+(1) The agent exhibits unstructured behavior, often repeating unnecessary actions.
 (2) It frequently toggles doors multiple times or picks up and drops keys without using them effectively.
@@
 (2) Upon encountering an obstacle, it either turns or moves in the opposite direction.
-Red-Blue Doors (1) The agent prioritizes opening the red door before the blue door, optimizing reward accumulation.
+Red–Blue
+Doors
+(1) The agent prioritizes opening the red door before the blue door, optimizing reward accumulation.
 (2) It successfully interacts with doors in a structured sequence, adhering to task constraints.
@@
 (3) In dense obstacle scenarios, occasional inefficiencies or unnecessary turns are observed.
-Table 2: Summarized insights on agent behaviors across different tasks.
+Table 3: Insights from agents’ SySLLM summaries in the MiniGrid environments.
+Agent Insights from Summaries
+Crafter
+Resource-collector
+(1) Strong Focus on Resource Collection: The agent consistently prioritizes gathering essential resources such as wood, stone, and drink, which are foundational for crafting tools and maintaining basic survival metrics.
+(2) Effective Basic Tool Crafting: Regular crafting of basic tools like wood pickaxes and swords enables the agent to enhance resource collection and engage in occasional combat.
+(3) Achievement Unlocking: The agent reliably unlocks achievements related to resource collection and basic tool crafting but struggles with more advanced achievements, highlighting a potential area for improvement.
+(4) Moderate Combat Engagement: The agent occasionally engages with zombies, using crafted tools for defense, showing moderate adaptability to threats but limited combat readiness overall.
+(5) Predictable Behavior: Episodes are characterized by high consistency in resource collection and basic tool crafting actions, with low variance across episodes.
+(6) Inconsistent Health and Exploration Management: While the agent effectively manages drink levels, it shows less consistency in food and health management and sacrifices exploration efficiency for achievement unlocking.
+Crafter
+Random
+(1) Ineffective Crafting and Resource Management: The agent frequently attempts crafting without the necessary resources or understanding of prerequisites, leading to repeated failures and minimal progress in achieving complex objectives.
+(2) Poor Survival Strategy: The agent consistently demonstrates ineffective survival behavior, including health depletion and poor management of food and drink levels, which hampers its ability to sustain itself in the game environment.
+(3) Limited Achievement Progression: While the agent reliably unlocks basic achievements like ‘wake_up’ and ‘collect_sapling’, it struggles to achieve more complex milestones that require crafting, resource management, or combat engagement.
+(4) Repetitive and Ineffective Actions: Episodes are marked by high frequencies of ‘noop’ actions and repetitive failed attempts at crafting, reflecting a lack of strategic adaptation and learning from past failures.
+(5) Lack of Combat Engagement: Lack of Combat Engagement: The agent shows minimal engagement with combat mechanics and fails to defend effectively against threats such as zombies and skeletons.
+(6) Predictable Behavior: Across episodes, the agent exhibits consistent, ineffective patterns of action, suggesting significant limitations in its decision-making processes and adaptability.
+Table 4: Insights from agents’ SySLLM summaries in the Crafter environment.
 6 Expert Evaluation of Summaries
-To validate the quality and alignment of SySLLM summaries, we conducted evaluations with RL experts.
+To evaluate the quality and alignment of SySLLM summaries, we conducted evaluations with RL experts.
 Specifically, six graduate students actively engaged in RL research with experience in training agents were recruited.
-These experts were divided into two groups: Experts 1, 2, and 3 evaluated the Minigrid-Unlock goal-directed and short-sighted agents, while Experts 4, 5, and 6 evaluated the remaining agents.
+These experts were divided into two groups: Experts 1, 2, and 3 evaluated the MiniGrid-Unlock goal-directed and short-sighted agents, while Experts 4, 5, and 6 evaluated the the rest of the MiniGrid agents.
+Experts 1,2, and 6 evaluated the agents from the Crafter environment.
 Each expert was shown videos approximately 120 seconds long, depicting the agents’ behaviors in the environment.
 Based on these videos, the experts were tasked with summarizing the agents’ behaviors.
-To ensure a fair comparison with SySLLM-generated summaries, we provided the experts with instructions that aligned closely with the prompts used for the language model (see Section F of the Appendix for detailed instructions).
+To ensure a fair comparison with SySLLM-generated summaries, we provided the experts with instructions that aligned closely with the prompts used for the language model (see G for detailed instructions).
 To evaluate how much of the experts’ observations were captured by the SySLLM summaries, we used a recall metric.
@@
 Each expert summary and the SySLLM summary were broken down into lists of key points, and a scoring system was applied: 1 point for a match with a SySLLM key point, 0.5 points for a partial match, and 0 points for no match.
-The recall scores, as shown in Table 3, ranged from 0.687 for the Unlock goal-directed agent, indicating moderate agreement, to 0.914 for the Crossing agent, reflecting a near-complete capture of expert insights.
-The overall mean recall score across all agents was 0.814, demonstrating substantial coverage of the experts’ points by the SySLLM summaries.
+The recall scores, as shown in Table 5, ranges from 0.687 for the Unlock goal-directed agent, indicating moderate agreement, to 0.914 for the Crossing agent, reflecting a near-complete capture of expert insights.
+The overall mean recall score across all agents is 0.840, demonstrating substantial coverage of the experts’ points by the SySLLM summaries.
 In the second evaluation phase, the experts were presented with key points from the SySLLM summaries that they had not mentioned in their own summaries.
@@
 Importantly, the precision calculation includes all SySLLM key points, encompassing both the points classified in this phase and those already evaluated during the recall assessment.
-The precision scores, as shown in Table 3, ranged from 0.769 (Dynamic Obstacles agent) to 0.864 (Short-Sighted agent), with an overall mean precision score of 0.819 across all agents.
+The precision scores, as shown in Table 5, range from 0.769 (Dynamic Obstacles agent) to 0.864 (Short-Sighted agent), with an overall mean precision score of 0.839 across all agents.
 In addition to quantifying correctness, this metric offers a way to evaluate potential hallucinations by identifying points that may not align with expert observations.
@@
 This point is classified as “not matched” by all experts, indicating a possible hallucination.
+To assess the reliability of expert annotations, we report both percentage agreement and Gwet’s AC1–metrics appropriate for imbalanced labeling tasks.
+The mean agreement reached 70%, and Gwet’s AC1 indicated substantial reliability (mean = 0.72), reflecting strong consistency across annotators.
+These results not only validate the quality of the annotations, but also provide important context for interpreting model performance: if human experts do not fully agree, perfect alignment between model and human labels is neither expected nor realistic.
+In this light, the model’s ability to approximate expert consensus is particularly notable.
 Taken together, the integration of recall and precision metrics demonstrates SySLLM’s capability to produce comprehensive and accurate summaries that closely align with expert observations.
 Agent Expert Recall Precision Mean Recall Mean Precision
-Unlock Goal-Directed 1 0.500 0.864 0.687 0.864
+Unlock
+Goal-Directed
+1 0.500 0.864 0.687 0.864
 2 0.643 0.864
 3 0.917 0.864
-Unlock Short-Sighted 1 0.800 0.846 0.878 0.839
+Unlock
+Short-Sighted
+1 0.800 0.846 0.878 0.839
 2 0.833 0.807
 3 1.000 0.923
-Dynamic Obstacles 4 0.583 0.692 0.739 0.769
+Dynamic
+Obstacles
+4 0.583 0.692 0.739 0.769
 5 0.833 0.692
 6 0.800 0.923
-Lava Gap 4 0.667 0.769 0.794 0.811
+Lava Gap
+4 0.667 0.769 0.794 0.811
 5 0.786 0.846
 6 0.929 0.818
-Red Blue Doors 4 0.857 0.767 0.871 0.834
+Red Blue
+Doors
+4 0.857 0.767 0.871 0.834
 5 0.857 0.867
 6 0.900 0.867
-Crossing 4 0.750 0.731 0.914 0.795
+Crossing
+4 0.750 0.731 0.914 0.795
 5 0.917 0.808
 6 1.000 0.846
+Crafter
+Resource-Collector
+7 0.938 0.893 0.931 0.871
+8 0.938 0.857
+9 0.917 0.864
+Crafter
+Random
+7 0.929 0.917 0.902 0.929
+8 1.000 0.958
+9 0.778 0.857
 Table 5: Recall and Precision of Expert Insights for SySLLM.
@@
 This variety allowed us to better assess how well the summaries captured different decision-making strategies and provided interpretable insights into the agents’ behavior.
-For completeness, we provide an intuitive explanation of HIGHLIGHTS in Section A of the Appendix.
+For completeness, we provide an intuitive explanation of HIGHLIGHTS in A.
 We assessed participants’ subjective preferences as well as their ability to identify agent behavior based on the summaries.
@@
 The metrics, adapted from Hoffman’s explanation satisfaction questionnaire [55], included Clarity, Understandability, Completeness, Satisfaction, Usefulness, Accuracy, Improvement, and Preference.
-These adaptations were tailored to the study’s specific needs and excluded irrelevant items (see Section F for the full scale).
+These adaptations were tailored to the study’s specific needs and excluded irrelevant items (see G for the full scale).
 Participants then evaluated the second summary, corresponding to the alternative condition (HIGHLIGHTS or SySLLM), using the same set of questions.
@@
 For $Q1$ ($\chi^{2}=1.0523$, $p=0.3050$, odds ratio (OR) = 0.5843), $Q2$ ($\chi^{2}=0.6703$, $p=0.4129$, OR = 0.7550), and $Q3$ ($\chi^{2}=0.7279$, $p=0.3936$, OR = 1.3875), the Chi-Square tests showed no statistically significant differences.
-Independent t-tests showed no statistically significant differences in confidence scores between the SySLLM and HIGHLIGHTS conditions for $Q1$ ($t=-0.6202$, $p=0.5359$) and $Q2$ ($t=0.5565$, $p=0.5785$).
+Independent t-tests showed no statistically significant differences in confidence scores between the SySLLM and HIGHLIGHTS conditions for $Q1$ (t=0.6202t=-0.6202, $p=0.5359$) and $Q2$ ($t=0.5565$, $p=0.5785$).
 However, for $Q3$, there was a statistically significant difference ($t=3.4197$, $p=0.0008$).
@@
 8 Discussion and Future Work
-We introduce SySLLM, leveraging LLMs to generate summaries of reinforcement learning agents.
-Its effectiveness is demonstrated across five distinct environments, including seven different agents.
-Expert evaluations show that SySLLM closely aligns with expert-generated summaries, capturing key aspects of agent behavior.
-User studies reveal that participants rate SySLLM summaries higher than HIGHLIGHTS on subjective metrics and perform equally well in agent identification tasks.
-These results provide evidence for the potential usefulness of the proposed approach.
-We now turn to discuss several limitations of the SySLLM method and possible ways to address them, as well as additional directions for future work.
-SySLLM relies on the availability of a captioning function, a widely accepted approach in language-integrated methodologies, as evidenced by prior research [13, 26, 20].
-In simulated environments, captioning is particularly straightforward, as ground truth states are readily accessible, and numerous RL environments, such as NetHack/MiniHack [32, 22] and text-based games [35, 11], already incorporate language-based features.
-In more complex or real-world scenarios, robust techniques such as object detection [41], cutting-edge captioning frameworks [36], action recognition systems [21], or advanced vision-language models [2] can be utilized.
-Alternatively, annotator models [13] or pretrained foundation models [37] offer viable solutions for generating captions.
-With continued advancements in general-purpose captioning models, we anticipate that readily available, off-the-shelf captioning tools will soon support an even broader range of tasks, further solidifying the practicality of this approach.
-Advancements in LLMs, particularly larger context windows, can extend this approach to large-scale and more complex environments, further validating its effectiveness and unlocking its potential to provide deeper insights into agent behavior.
-Future work could explore a question-answering (QA) framework, allowing users to query the agent’s policy directly.
-For example, users could ask why specific actions were taken or predict the agent’s behavior in new scenarios.
-This interactive system would provide deeper insights into the agent’s decision-making process, offering a flexible and intuitive way to understand its behavior.
-Another avenue for future work is generating multimodal summaries, combining natural language explanations with visual representations of key moments, trajectories, and decisions.
-That is, for each pattern identified by the LLM, the summary would also include demonstrations of the pattern extracted from agent traces.
-This approach could enhance user understanding by integrating text and visuals to more effectively convey complex policies.
-Visual elements from the agent’s experience dataset might help mitigate hallucinations by grounding explanations in observable behavior, potentially improving reliability and trustworthiness.
+In this paper, we introduced SySLLM, an innovative framework leveraging large language models to synthesize reinforcement learning policies into clear, structured narratives.
+Our method addresses key shortcomings of traditional extractive-visual methods by moving away from fragmented, scenario-specific snapshots toward holistic, language-based summaries.
+This shift significantly reduces the effort required for users to interpret complex agent behaviors, facilitating better comprehension of strategic decisions and adaptability.
+Our evaluations provide strong evidence that SySLLM effectively captures nuanced policy behaviors and aligns closely with expert-generated summaries.
+The qualitative alignment with human experts highlights the capability of language models to accurately convey intricate behavioral patterns and strategic nuances, reinforcing the effectiveness of a textual-abstractive approach.
+Furthermore, user preference studies demonstrate clear advantages in interpretability and user engagement over visual-based summarization methods.
+The broader implications of SySLLM suggest promising new directions for research and practice.
+We advocate for increased exploration into abstractive-textual explanations in global explanation in reinforcement learning, encouraging the community to investigate advanced techniques such as fine-tuning, in-context learning, and integration with reasoning frameworks.
+Embracing this paradigm can lead to significant advancements in how we communicate complex policies, ultimately enhancing trust, transparency, and adoption of reinforcement learning in critical real-world scenarios.
+Captioners
+Our approach relies on the availability of effective captioning functions, a common requirement in language-integrated methodologies as seen in prior research [35, 46, 45].
+In simulated environments, this dependency is relatively straightforward to address, as ground truth states are accessible, and many RL environments—such as NetHack/MiniHack [56, 57] and text-based games [58, 59]—already incorporate language-based features.
+For more complex or real-world scenarios, however, implementations would require robust object detection [60], advanced captioning frameworks [61], action recognition systems [62], or vision-language models [63].
+Alternatively, annotator models [35] or pretrained foundation models [64] could generate captions.
+The continued advancement of general-purpose captioning models suggests that off-the-shelf solutions will soon support an even broader range of tasks, enhancing the practicality of our approach.
+LLMs Limitations
+The current performance of SySLLM is bounded by the capabilities and limitations of the underlying language models.
+Context window constraints restrict the amount of agent trajectory data that can be analyzed at once, potentially limiting the ability to identify long-term patterns or rare behaviors.
+Additionally, LLMs can occasionally generate plausible-sounding but incorrect explanations (hallucinations), particularly when faced with unusual or complex agent behaviors.
+As LLM technology evolves, particularly with expanded context windows and improved factuality, we anticipate that these limitations will diminish, enabling more comprehensive and reliable agent behavior summaries.
+Building on SySLLM’s strengths while addressing its limitations, we identify several promising directions for future research:
+Interactive Question-Answering Framework
+A promising avenue for future work is extending SySLLM to support an interactive question-answering (QA) framework.
+This would enable users to directly query the agent’s behavior, asking targeted questions such as “Why did the agent choose action A over B in this situation?” or “How would the agent respond to condition X?” Such interactivity could transform SySLLM from a static summarization tool into a dynamic interface for policy exploration, offering a more flexible and intuitive means of understanding agent behavior.
+Prior work on logical querying of agents, such as [65], highlights the feasibility and potential of QA-based explanations in RL.
+Inspired by these developments, we envision a natural language-based QA interface grounded in narrative trajectories, allowing users to probe specific aspects of agent decision-making in a more accessible way.
+Multi-modal Summaries
+Integrating natural language explanations with visual representations would significantly enhance the communicative power of SySLLM.
+For each behavioral pattern identified by the language model, the system could provide corresponding visual evidence extracted from agent traces such as state transitions, action sequences, or critical decision points.
+This multi-modal approach would ground abstract explanations in concrete examples, potentially mitigating hallucination issues while making complex policies more accessible to users with diverse backgrounds and expertise levels.
+Comparative Analysis Framework
+Future work could extend SySLLM to automatically compare multiple agents’ behaviors, identifying similarities, differences, and relative strengths.
+This comparative capability would be particularly valuable for evaluating alternative training approaches, understanding the effects of hyperparameter changes, or analyzing how agent behavior evolves during training.
+By articulating these comparisons in natural language, SySLLM could provide intuitive insights into the factors that influence agent performance and behavior.
 Acknowledgments
@@
 The agent’s performance could potentially be enhanced by expanding its field of vision or optimizing its pathfinding algorithms to increase consistency in achieving higher rewards in fewer steps.
-Appendix E Implementation Details
+Appendix E Crafter Summarization Prompts
+You are analyzing multiple episode summaries of an agent’s behavior in the Crafter environment.
+Your goal is to identify consistent behavioral patterns across episodes and develop an objective behavioral profile.
+[Cross-Episode Analysis Framework]
+Systematically evaluate these key dimensions:
+Behavioral Consistency
+- Identify actions and strategies that appear consistently across episodes
+- Calculate the variance in action distributions between episodes
+- Note any evolution or change in behavior across sequential episodes
+- Determine if the agent displays consistent preferences or purely situational responses
+Achievement Patterns
+- Calculate achievement unlock rate and consistency across episodes
+- Identify which achievements are most frequently obtained
+- Analyze the typical sequence or prerequisites leading to achievements
+- Assess whether achievement patterns suggest intentional pursuit or incidental acquisition
+Resource Priorities
+- Identify primary resources consistently targeted across episodes
+- Analyze typical crafting sequences when resources are available
+- Evaluate how consistently the agent manages inventory
+- Determine if there are clear resource collection preferences
+Environmental Interaction Patterns
+- How consistently does the agent navigate the environment?
+- Identify common responses to specific environmental features
+- Analyze patterns in exploration vs. exploitation behavior
+- Evaluate adaptation to threats, opportunities, and constraints
+Decision-Making Characteristics
+- Identify the apparent decision criteria for different action choices
+- Analyze how the agent balances short-term vs. long-term needs
+- Evaluate how predictable the agent’s responses are to similar situations
+- Assess whether actions appear purposeful or random
+[Output Instructions]
+1. Begin with a "Behavioral Profile" summarizing the agent’s most consistent traits
+2. Include a "Statistical Analysis" section with quantitative breakdowns of action patterns
+3. Provide a "Decision Pattern Analysis" detailing how the agent makes choices
+4. Add an "Achievement Analysis" showing typical patterns in achievement progression
+5. Conclude with "Behavioral Consistency Assessment" that evaluates how predictable the agent is
+Give the agent a label based on its observed behavior and justify your choice.
+Your analysis should be based entirely on observable patterns.
+If the agent shows highly inconsistent behavior across episodes, explicitly detail this with supporting evidence.
+Focus on describing what the agent does consistently, rather than speculating on why it might do so.
+Appendix F Implementation Details
+F.1 MiniGrid
 We employed the PPO algorithm from the stable-baselines3 library for our policy network, which takes as input a $K\times K\times 3$ encoded image and a mission string, the latter being encoded using a one-hot scheme.
@@
 Entropy Coefficient 0.001 0.001 0.01 0.001 0.001 0.01
-Learning Rate $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$
+Learning Rate $\times 10^{-4} 1×1$$\times 10^{-4}$ $\times 10^{-4} 1×1$$\times 10^{-4}$ $\times 10^{-4} 1×1$$\times 10^{-4}$
 Clip Range 0.2 0.2 0.2 0.2 0.2 0.2
 Table 8: Hyper-parameters for the PPO algorithm applied to all six agents.
-Appendix F Experts instructions:
+F.2 Crafter
+We implemented DreamerV3 for our agent, using a state-of-the-art world model-based reinforcement learning approach.
+The agent processes $64\times 64\times 3$ RGB observations from the Crafter environment.
+The world model consists of three key components: an encoder network, a recurrent state-space model (RSSM), and a decoder network.
+The encoder transforms raw pixel observations into a 1024-dimensional embedding space using a convolutional neural network with a depth of 96 channels.
+The RSSM, which forms the core of the agent’s predictive capabilities, utilizes a deterministic state of dimension 4096 and a stochastic state represented as a 32-dimensional random variable, allowing the agent to account for environment stochasticity.
+For temporal dynamics, we employed a GRU cell with 1024 hidden units.
+The decoder reconstructs observations using transposed convolutions, enabling the model to learn compact state representations through reconstruction loss.
+For policy learning, we used an actor-critic architecture with 5-layer MLPs for both actor and critic, where the actor employs a categorical distribution over the 17 discrete actions available in Crafter.
+The agent was trained using the ”reinforce” gradient strategy for imagination-based policy optimization, with a $\lambda$-return horizon of 15 steps and a discount factor of 0.997.
+Training was conducted for $10^{6}$ environment steps using 8 parallel environments, with a batch size of 32 and sequence length of 64.
+We employed a model learning rate of $10^{-4}$ and an actor learning rate of $3\times 10^{-5}$, optimized using Adam.
+Appendix G Experts instructions:
 General Instructions:
@@
 The agent description should be at least 100 words. Provide approximately 5 key insights.
-Appendix G Scale Used in Task 1
+Appendix H Scale Used in Task 1
 In Task 1 of our study, we utilized a 7-point Likert scale to evaluate participants’ perceptions and understanding of the agent’s behavior as presented in both the video summaries and the natural language summaries.
@@
 This scale aimed to capture various dimensions of participant satisfaction and understanding, contributing to the overall evaluation of the summaries’ utility in the context of our research.
-Appendix H Systematic Exploration of the Prompt Design
+Appendix I Systematic Exploration of the Prompt Design
 The creation of the final prompt was achieved through a structured and iterative exploration process.

### v3

@@
 From Actions to Words: Towards Abstractive-Textual Policy Summarization in RL
-Abstract
-Policies generated by Reinforcement Learning (RL) algorithms are difficult to explain to users, as they emerge from the interaction of complex reward structures and neural network representations.
-Consequently, analyzing and predicting agent behavior can be challenging, undermining user trust in real-world applications.
-To facilitate user understanding, current methods for global policy summarization typically rely on videos that demonstrate agent behavior in a subset of world states.
-However, users can only watch a limited number of demonstrations, constraining their understanding.
-Moreover, these methods place the burden of interpretation on users by presenting raw behaviors rather than synthesizing them into coherent patterns.
-To resolve these issues, we introduce SySLLM (Synthesized Summary using Large Language Models), advocating for a new paradigm of abstractive-textual policy explanations.
-By leveraging Large Language Models (LLMs)—which possess extensive world knowledge and pattern synthesis capabilities—SySLLM generates textual summaries that provide structured and comprehensible explanations of agent policies.
-SySLLM demonstrates that LLMs can interpret spatio-temporally structured descriptions of state-action trajectories from an RL agent and generate valuable policy insights in a zero-shot setting, without any prior knowledge or fine-tuning.
-Our evaluation shows that SySLLM captures key insights, such as goal preferences and exploration strategies, that were also identified by human experts.
-Furthermore, in a large-scale user study (with 200 participants), SySLLM summaries were preferred over demonstration-based summaries (HIGHLIGHTS) by a clear majority (75.5%) of participants.
-journal: Artificial Intelligence
+Abstract.
+Explaining reinforcement learning agents is challenging because policies emerge from complex reward structures and neural representations that are difficult for humans to interpret.
+Existing approaches often rely on curated demonstrations that expose local behaviors but provide limited insight into an agent’s global strategy, leaving users to infer intent from raw observations.
+We propose SySLLM (Synthesized Summary using Large Language Models), a framework that reframes policy interpretation as a language-generation problem.
+Instead of visual demonstrations, SySLLM converts spatiotemporal trajectories into structured text and prompts an LLM to generate coherent summaries describing the agent’s goals, exploration style, and decision patterns.
+SySLLM scales to long-horizon, semantically rich environments without task-specific fine-tuning, leveraging LLM world knowledge and compositional reasoning to capture latent behavioral structure across policies.
+Expert evaluations show strong alignment with human analyses, and a large-scale user study found that 75.5% of participants preferred SySLLM summaries over state-of-the-art demonstration-based explanations.
+Together, these results position abstractive textual summarization as a paradigm for interpreting complex RL behavior.
+[^1] Code available at https://github.com/saharad1/SySLLM
+Key words and phrases:
+Reinforcement Learning, Policy Summarization, Explainable AI
+{CCSXML}
+¡ccs2012¿ ¡concept¿ ¡concept_id¿10010147.10010178.10010179.10010182¡/concept_id¿ ¡concept_desc¿Computing methodologies Natural language generation¡/concept_desc¿ ¡concept_significance¿500¡/concept_significance¿ ¡/concept¿ ¡concept¿ ¡concept_id¿10010147.10010257.10010258.10010261¡/concept_id¿ ¡concept_desc¿Computing methodologies Reinforcement learning¡/concept_desc¿ ¡concept_significance¿500¡/concept_significance¿ ¡/concept¿ ¡/ccs2012¿
+1. Introduction
+Reinforcement learning (RL) agents are increasingly deployed in sequential decision-making domains, yet their policies remain opaque to human stakeholders.
+This opacity limits trust, adoption, and effective debugging.
+While many explainable reinforcement learning (XRL) methods provide local insight into individual states or decisions, our focus is on global explanations that aim to capture an agent’s strategy across trajectories.
+Existing paradigms face trade-offs between expressiveness, scalability, and faithfulness.
+Extractive methods such as saliency maps greydanus2018visualizing; puri2019explain; atrey2019exploratory; samadi2024safe and demonstration-based policy summaries amir2018agent; frost2022explaining; liu2023learning; deshmukh2023explaining capture only fragments of behavior.
+Symbolic approaches such as rules or decision trees sequeira2020interestingness; peng2022inherently; mccarthy2022boolean; parham2023explaining often collapse in high-dimensional or partially observed environments.
+Across these approaches, a core flaw persists: they transform observations but do not synthesize them into holistic accounts, leaving users to reconstruct intent, adaptability, and failure modes from fragmented evidence anderson2020mental; amitaisurvey.
+Large language models (LLMs) appear to offer a promising alternative.
+Their capacity for abstraction, compositional reasoning, and natural language generation zhao2023survey suggests that they could distill trajectories into human-readable accounts of agent behavior.
+However, applying them to RL policies is far from straightforward.
+LLMs are trained on static text-based corpora, while RL agents generate dynamic spatio-temporal trajectories grounded in states, actions, and rewards li2024advancing.
+This mismatch in modality makes faithful policy summarization non-trivial.
+Specifically, naive application risks producing fluent and plausibly sounding behavior descriptions that do not accurately reflect the underlying policy.
+To address this challenge, we propose a textual-abstractive paradigm for policy summarization, in which explanations are expressed as natural language narratives integrating evidence from a database of policy execution traces, which we term the textual experience buffer (TEB).
+Unlike extractive or symbolic methods, this paradigm explicitly targets policy-level regularities and behavioral motifs.
+We formalize the task as a mapping from an agent’s experience buffer to a textual summary, guided by the desiderata of expressiveness, scalability, and faithfulness.
+To realize these principles in practice, we adopt a conceptual optimization view in which summaries balance coverage, parsimony, and fidelity.
+This formulation grounds both our methodology and evaluation, ensuring that system design choices are principled and directly address the unique challenges of applying LLMs to RL.
+Building on this formulation, we introduce SySLLM, a framework that leverages LLMs to generate global policy summaries.
+SySLLM operates in two stages: agent–environment trajectories are first transformed into structured natural language descriptions of observations and actions, which are then synthesized into higher-level accounts through carefully designed prompting.
+To scale beyond context window limitations, SySLLM performs hierarchical summarization over large buffers.
+To mitigate variability in LLM outputs, it generates multiple candidate summaries and aggregates them into a consensus using embedding-based similarity.
+Together, these mechanisms ensure that SySLLM produces summaries that are general enough to capture policy-level regularities while specific enough to reflect distinctive behaviors.
+We evaluated SySLLM across five MiniGrid environments and the Crafter domain, covering nine qualitatively distinct agent policies.
+Expert evaluation shows a strong alignment between SySLLM summaries and expert summaries, achieving high recall and precision scores that reflect faithful coverage of expert-identified behaviors.
+A user study with 192 participants further demonstrates that users strongly prefer textual summaries over demonstration-based summaries such as HIGHLIGHTS-DIV amir2018highlights, while performing equally well or better on policy identification tasks.
+Our main contributions are threefold: (1) we introduce and formalize the task of abstractive-textual policy summarization in RL, framing it as a mapping from an agent’s experience buffer to natural language narratives defined by desiderata of expressiveness, scalability, and faithfulness; (2) we present SySLLM, a framework that leverages LLMs to synthesize structured trajectory descriptions into coherent global summaries, incorporating hierarchical summarization and consensus aggregation to address long horizons and variability in outputs, and (3) we provide extensive empirical validation through expert evaluations and a large-scale user study, demonstrating that SySLLM produces faithful summaries that are strongly preferred by participants over state-of-the-art demonstration-based baselines.
+2. Related Work
+Prior work in XRL spans both local explanations of individual decisions and global summaries of agent behavior, with the shared goal of improving policy interpretability for humans milani2022survey.
+Saliency and visualization methods highlight influential inputs greydanus2018visualizing; huber2022benchmarking but are often local and fragile; demonstration-based methods summarize behavior via selected trajectories amir2018agent; deshmukh2023explaining but place interpretive burden on users; surrogate models distill policies into rules or decision trees bastani2018verifiable; coppens2019distilling, though with fidelity–scalability trade-offs; and causal or reward-based explanations madumal2020explainable; juozapaitis2019explainable provide structured insights but typically require access to internals or domain expertise.
+Collectively, these techniques emphasize fragments or simplified proxies rather than synthesizing global accounts of an agent’s strategy.
+Large language models (LLMs) bring complementary capabilities of abstraction, reasoning, and fluent text generation brown2020language.
+While prior work mainly uses LLMs during training—for example, guiding exploration or constructing world models du2023guiding; guan2023leveraging—their potential as explanation generators has received less attention.
+Some studies prompt LLMs to narrate behavior in real time wang2023voyager; yao2023react or to build symbolic simulators ahn2022can, but these efforts are often ungrounded in actual dynamics and focus on local rather than global explanations.
+Natural language explanations for RL agents have also been studied.
+The early template-based methods hayes2017improving; peng2022inherently prioritized accessibility, but were brittle, while neural rationalization approaches translated trajectories into free form text ehsan2018rationalization; mccalmon2022caps.
+These provide interpretability but generally rely on handcrafted structures, focus on local justifications, or lack scalability across diverse scenarios.
+They rarely capture holistic behavioral patterns or leverage broader common sense knowledge.
+These limitations motivate our textual-abstractive paradigm, where LLMs synthesize coherent global summaries of agent behavior.
 […]
+Diagram of the data collection pipeline.
+The environment produces an observation grid showing objects like a yellow key and door.
+The observation is converted into a text description, then passed to an experience dataset.
+The agent’s policy network takes the observation and outputs an action, such as moving forward.
+The action and observation captions are stored in the dataset along with rewards.
+Figure 1. Collecting the textual experience buffer (Section 4.2).
+3. Problem Formulation
+In this section, we formalize the abstractive-summarization problem.
+Specifically, we describe the setting, define the summarization task, and introduce conceptual principles that guide both our methodology and our evaluation.
+3.1. Setting
+We consider an RL environment modeled as a partially observable Markov decision process (POMDP):
+$\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{O},T,O,R,\gamma\rangle,$
+where $\mathcal{S}$ is the state space, $\mathcal{A}$ the action space, $\mathcal{O}$ the observation space, $T(s^{\prime}\mid s,a)$ the transition kernel, $O(o\mid s,a)$ the observation function, $R(s,a)$ the reward function, and $\gamma\in(0,1]$ the discount factor.
+An agent follows a stochastic policy
+$\pi:\mathcal{O}\to\Delta(\mathcal{A}),$
+which maps each observation $o\in\mathcal{O}$ to a distribution over actions. The interaction induces a distribution over trajectories
+$\tau=(o_{0},a_{0},r_{0},\dots,o_{T}),\quad\tau\sim\pi,T,O.$
+For summarization, we assume access not only to isolated trajectories but to an experience buffer
+$\mathcal{B}_{\pi}=\{\tau_{1},\dots,\tau_{N}\},$
+which aggregates multiple episodes sampled from $\pi$. This buffer serves as the raw material based on which textual explanations are generated.
+3.2. Policy Summarization Task
+We define a policy summarizer as a mapping
+$f:\mathcal{B}_{\pi}\to\mathcal{T},$
+where $\mathcal{T}$ denotes the space of abstractive textual explanations.
+We require $f$ to approximate three key principles.
+Expressiveness: capture recurring behavioral patterns (e.g.
+“the agent prioritizes unlocking doors before exploring rooms”) rather than isolated actions.
+Scalability: operate over long horizons and large buffers while maintaining concise summaries.
+Faithfulness: reflect the actual distribution of behaviors under $\pi$, avoiding hallucinated or spurious strategies.
+Conceptual Optimization View
+We frame summarization as a conceptual optimization problem:
+$T^{*}=\arg\max_{T\in\mathcal{T}}\;\mathcal{U}(T\mid\mathcal{B}_{\pi}),$
+where $\mathcal{U}$ is a utility function balancing: coverage (operationalizing expressiveness by accounting for recurring behaviors), parsimony (operationalizing scalability through concise abstraction), and fidelity (operationalizing faithfulness via alignment with empirical evidence in $\mathcal{B}_{\pi}$).
+This formulation is not solved directly.
+Instead, it provides a guiding lens for both methodology and evaluation: SySLLM instantiates coverage through multi-trajectory aggregation and hierarchical summarization, parsimony through representative summary selection, and fidelity through expert alignment and user validation.
+In the following sections, we detail how these principles are operationalized in practice and evaluated empirically.
 […]
-1 Introduction
-Reinforcement learning (RL) agents are increasingly deployed in high-stakes domains, yet their opaque decision-making processes remain a barrier to trust and adoption.
-Explainable RL (XRL) methods aim to bridge this gap by surfacing interpretable accounts of agent behavior.
-Despite this promise, current XRL methods face significant limitations.
-Saliency maps [1, 2, 3, 4] highlight critical states but offload pattern discovery to users.
-Trajectory demonstrations [5, 6, 7, 8] constrain the understanding to cherry-picked scenarios.
-Rule extraction methods [9, 10, 11, 12], while providing structured representations, often struggle to capture the full complexity of learned policies.
-Many of these approaches share a common challenge: they present transformed observations rather than synthesized insights, requiring users to manually extrapolate agent intent, adaptability, and failure modes from fragmented evidence [13].
-Recent advances in large language models (LLMs) offer a transformative alternative.
-With their capacity for abstraction, commonsense reasoning, and natural language generation [14], LLMs hold unique potential to synthesize agent trajectories into human-interpretable narratives.
-However, this promise comes with a challenge, LLMs’ static text-based training contrasts with the dynamic, spatiotemporal nature of reinforcement learning environments [15].
-Early applications of LLMs in RL, such as world modeling [16], hint at their promise, but generating faithful global policy summaries—cohesive accounts that generalize beyond isolated states—remains unexplored.
-Figure 1: Collecting the experience dataset (Section 4.3).
-In this work, we advocate for an abstractive-textual approach for policy summarization.
-We introduce SySLLM, a framework that leverages LLMs to generate holistic summaries of RL policies by systematically distilling agent trajectories into structured narratives.
-SySLLM addresses the limitations of previous methods through a two-stage process: (1) converting raw agent-environment interactions into temporally grounded language descriptions (Fig.
-1), and (2) prompting LLMs to synthesize these into coherent summaries that capture behavioral patterns, strategic tendencies, and environment-specific adaptations (Fig.
-2).
-SySLLM performs impressively well—it captures agent preferences (as induced by different reward functions), identifies sub-optimal behavior in under-trained agents, and distinguishes between diverse exploration strategies.
-Beyond summarizing the agent’s behavior, SySLLM offers insights into the effectiveness of the agent’s strategy.
-SySLLM demonstrates that LLMs can interpret spatiotemporally structured descriptions of state-action trajectories from an RL agent.
-While these descriptions are written in natural language, they differ significantly from the kinds of text LLMs are typically trained on, such as stories or dialogue.
-Instead, they represent a distinct modality: sequential, decision-driven behavior grounded in the dynamics of an environment.
-Despite this, SySLLM shows that LLMs can process such unfamiliar input and, in a zero-shot setting—without any prior knowledge or fine-tuning—generate valuable policy insights.
-Our framework is validated through multi-faceted evaluation.
-First, comparisons with RL experts reveal that SySLLM summaries align closely with expert annotations, capturing nuanced behaviors that demonstration-based methods do not explicitly highlight.
-Second, a large-scale user study shows that participants prefer SySLLM over state-of-the-art visual summaries such as HIGHLIGHTS [17], while also achieving equal or better accuracy in agent identification tasks.
-In summary, our main contributions are the following:
-  • We introduce SySLLM, a framework that takes a step toward generating abstractive-textual explanations of RL policies using LLMs.
-SySLLM synthesizes structured narratives that reveal agent intent, adaptability, and limitations by systematically analyzing agent trajectories, enabling users to grasp these behaviors without manual pattern matching.
-SySLLM eases the burden of interpreting agent behavior by digesting complex dynamics into coherent summaries.
-To the best of our knowledge, this is the first work to produce global, abstractive explanations of an RL agent’s policy in natural language.
-  • Generated summaries align with RL experts.
-We evaluate the effectiveness of SySLLM by comparing its summaries with those written by RL experts, demonstrating substantial alignment and accurate coverage of key policy details.
-To evaluate faithfulness—an area often prone to issues in automatic summarization [18, 19]—we additionally compute precision scores.
-Our results show that SySLLM maintains high precision overall, indicating that its summaries rarely introduce unsupported claims.
-  • Users prefer our textual summaries.
-We conduct a large-scale user study with 200 participants showing that participants clearly prefer SySLLM over a leading demonstration-based policy summarization method on a summary preference task (75.5%), while performing equally well or better in an agent identification task.
-Figure 2: Generating natural language global policy summaries (Section 4.4).
-2 Background and Related Work
-2.1 Explainable Reinforcement Learning
-Explainable reinforcement learning (XRL) aims to make the behavior of RL agents more understandable to human users.
-Traditional XRL techniques fall into several major categories: saliency-based methods, demonstration-based summaries, surrogate models, and causal explanations [20].
-Saliency and Visualization Techniques
-Early XRL work often adopted visualization tools such as saliency maps [1, 2, 3, 4] to indicate which parts of the input space influenced the agent’s decisions.
-While these tools provide some insight, they require users to manually interpret the spatial and temporal relevance of highlighted regions, leading to potential misinterpretation.
-Studies show these methods are local and fragile [21], providing little insight into the long-term policy structure.
-Demonstration-Based Summaries
-Another common approach involves selecting representative episodes or critical states that demonstrate agent behavior [5, 6, 7, 8].
-These methods allow users to observe behavior directly, but rely heavily on cherry-picked scenarios, with high cognitive load placed on the user to infer strategic intent from limited data.
-Work by [22] shows that users often struggle to generalize from such demonstrations, particularly when the agent behavior is stochastic.
-Surrogate Models
-An alternative strategy uses model distillation [23] to learn interpretable approximations of agent behavior, such as decision trees [24, 25].
-These methods strike a balance between interpretability and fidelity: overly simple models fail to capture nuances of the agent’s strategy, while complex surrogates become as opaque as the original policy.
-Recent works explore bounding MDPs [26] or abstract policy graphs [27] to improve this trade-off, but scalability remains an issue.
-Causal Explanations
-A complementary direction models the agent’s reasoning through structural causal models or influence graphs [28, 29].
-These allow users to query “why” an agent took a specific action and offer contrastive explanations that align with how humans seek understanding.
-However, such models often require significant prior knowledge of the domain or agent internals, limiting their applicability.
-Reward Decomposition
-Other methods explain agent behavior by breaking down the overall reward into meaningful components, such as safety, efficiency, or task-specific objectives.
-This allows explanations that clarify how each reward type influences the agent’s decisions and highlights trade-offs between competing objectives [30, 31, 32].
-These approaches help users understand agent priorities and decision rationale by making the contribution of each reward component explicit.
-2.2 Large Language Models in Reinforcement Learning
-Pre-trained Large Language Models (LLMs) have revolutionized NLP through their ability to distill patterns from vast datasets into coherent text.
-Autoregressive models like GPT [33] generate human-like narratives by predicting token sequences, while encoder-decoder architectures excel at summarizing complex inputs [34].
-Those recent advances have introduced new possibilities for interacting with and explaining RL agents.
-While much prior work integrates LLMs into the training process e.g., guiding exploration [35] or building world models [16], less attention has been paid to their potential as explanation generators.
-[36] explore using LLMs to generate local post-hoc explanations of individual decisions based on distilled behavior representations, but such approaches do not aim to provide global summaries of agent behavior.
-Trajectory Annotation and Feedback
-Some works use LLMs to narrate agent behavior in real-time or provide textual feedback.
-[37, 38] demonstrate that LLMs can be prompted to interpret actions or suggest improvements mid-episode.
-These studies show promise for LLM-based commentary, though they often lack grounding in actual agent dynamics, leading to inconsistencies.
-Model-Based Reasoning and Simulation
-In world modeling applications, LLMs have been used to generate or refine symbolic models of environment dynamics [39].
-Here, the LLM acts as a high-level planner or simulator, encoding causal knowledge that guides behavior.
-This further validates the hypothesis that LLMs can represent abstract behavioral structures, which is essential for high-level summarization.
-These applications largely treat LLMs as agents or environment models, not as explainers.
-When LLMs are used for explanation, they must be carefully grounded in observed experience to avoid plausible-sounding but inaccurate generalizations [40].
-2.3 Natural Language Explanations for RL Agents
-The use of natural language as a modality for RL explanations offers distinct advantages over visual or rule-based approaches.
-Language allows for abstraction, temporal coherence, and commonsense reasoning—all of which are challenging to express in saliency maps or decision trees.
-Template-Based Explanations
-Early works, such as [41], demonstrated that combining learned MDP structures with templated natural language generation could enable contrastive explanations (e.g., “Why didn’t you do X?”) through predicate-based state abstraction, allowing generalization across policy architectures.
-Later, simpler approaches reemerged, such as [10], which employed rigid, static templates to directly translate policies into interpretable statements (e.g., “If near wall, turn left”), prioritizing accessibility but inheriting brittleness and limited expressivity.
-More recent systems like [42] also rely on predefined action-condition mappings, facing similar limitations.
-Rationalization and Free-Form Generation
-Subsequent work adopted neural methods to generate textual rationales.
-[43] introduced AI rationalization training sequence-to-sequence models to translate trajectories into natural language.
-Later, [44] introduced graph-based policy summaries that enhance user understanding by clustering states and abstracting the state space.
-While those natural language explanations methods offer interpretability, they are inherently limited by their reliance on handcrafted rules and rigid structures, making them brittle and unable to express nuanced or unexpected behaviors.
-Rationalization methods, which translate individual trajectories into free-form text, tend to focus on local justifications rather than providing a holistic view of agent behavior.
-These approaches often lack scalability across diverse scenarios and struggle to generalize patterns observed over many episodes.
-Moreover, they typically do not leverage broader world knowledge or incorporate high-level reasoning, which can limit the clarity, depth, and utility of the resulting explanations for end users.
-3 Motivation for Abstractive-Textual Explanations
-Despite recent advances in explainable reinforcement learning (XRL), existing methods face a persistent trade-off between expressiveness, scalability, and accessibility.
-Extractive approaches—such as saliency maps or curated demonstrations—highlight fragments of agent behavior but often fail to integrate them into coherent, high-level narratives.
-Symbolic models, including logical rules and decision trees, offer structural clarity but require handcrafted features and often struggle to scale in complex or high-dimensional settings.
-These limitations point to a need for explanation methods that can synthesize global behavioral insights—capturing patterns, adaptation, and strategy over time—without relying on structured input representations or human-crafted abstractions.
-We advocate for a textual-abstractive paradigm in XRL, where natural language is used to summarize agent behavior across episodes.
-Text offers a flexible medium for conveying temporal dependencies, recurring behaviors, and decision-making logic in a format accessible to non-experts.
-Unlike extractive or symbolic methods, textual explanations can operate over raw behavioral data without predefined rules or model internals, making them broadly applicable across domains.
-Recent advances in large language models (LLMs) open new opportunities for explainability in reinforcement learning.
-With SySLLM, we introduce a textual-abstractive paradigm that uses natural language to summarize agent behavior over time—capturing patterns, strategic intent, and adaptation.
-In this work, we show that LLMs are particularly well-suited for this approach, when guided by an appropriate framework, they can recognize sequential structure and express it in fluent, context-aware language.
-We view this paradigm as a holistic and scalable complement to existing XRL methods, and as part of a broader shift toward language-based interfaces for interpreting learned behavior.
-4 SySLLM Method
-Our SySLLM (Synthesized Summary using LLMs) method consists of several key components, which we formalize in Algorithm 1.
-In the first phase (illustrated in Figure 1), we run the agent in its environment and perform a captioning procedure to convert experiences into natural language.
-These captioned experiences are logged into an experience dataset (Sections 4.2 and 4.3).
-In the second phase, which we illustrate in Figure 2, we take the experience dataset as input and output a natural language summary that aims to capture patterns in the agent’s policy (Section 4.4).
-To do so, we compose a prompt with the formatted experience dataset that is then sent to an LLM to generate the summary.
-4.1 Problem Statement
-We define a partially observed Markov Decision Process (POMDP) as a tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{O},T,O,R,\gamma\rangle$, where observations $o\in\mathcal{O}$ are derived from states $s\in\mathcal{S}$ and actions $a\in\mathcal{A}$ via $O(o|s,a)$.
-The dynamics are governed by the transition function $T(s^{\prime}|s,a)$, with $R$ as the reward function and $\gamma$ as the discount factor.
-POMDPs generalize Markov Decision Processes (MDPs), where observations fully reveal the state.
-The agent operates under a stochastic policy $\pi:\mathcal{O}\rightarrow\Delta(\mathcal{A})$, mapping each observation $o\in\mathcal{O}$ to a probability distribution over actions $\mathcal{A}$.
-The objective of the policy is to maximize the expected cumulative discounted reward.
-Given a POMDP environment and the agent’s policy, our goal is to generate a natural language summary that encapsulates the agent’s strategy and decision-making process.
-This summary explains the policy’s operational dynamics, effectiveness, and behavior across scenarios.
-Formally, a summarization is a mapping from $(\mathcal{M},\Pi)$ to $\mathcal{T}$, where $(\mathcal{M},\Pi)$ represents environment-policy pairs, and $\mathcal{T}$ is the set of natural language summaries.
-4.2 Captioner
-As will be detailed in Section 4.3, our approach utilizes captioners to leverage LLMs for summarizing an agent’s policy.
-We convert each observation $o\in\mathcal{O}$ and action $a\in\mathcal{A}$ into natural language descriptions that highlight their key features.
-In this subsection, we formalize the role and function of the captioners.
-Observation Captioner:
-The observation captioner, $C_{\text{observation}}$, translates observations into natural language descriptions:
-$C_{\text{observation}}:\mathcal{O}\rightarrow\Sigma^{\star}.$
-Where $\Sigma^{\star}$ represents the set of all possible strings.
-This function distills complex state information into comprehensible text, such as translating a grid map into a description of object locations.
-Action Captioner:
-Similarly, the action captioner, $C_{\text{action}}$, maps actions to natural language representations:
-$C_{\text{action}}:\mathcal{A}\rightarrow\Sigma^{\star}.$
-Similar to past work [45, 46] we assume the availability of a captioner language annotator.
-The process of generating captions is supported by a range of established methods, which are further discussed in Section 8.
-4.3 Collecting the Agent’s Experience Dataset
-We describe the first phase of Algorithm 1, detailed in Lines 114.
-We assume that the environment ($ENV$) supports two functions.
-The first function is RESET, which resets the environment to its initial state and returns the initial observation, and STEP, which takes an action as input and returns the next observation, the reward for that action, a boolean indicating whether the episode has ended, and additional information.
-First, we initialize the experience dataset $ED$ (Line 1).
-The for loop (Line 2) iterates over multiple episodes, initializing the step counter $t=1$ at the beginning of each episode (Line 3) and obtaining the initial observation $o_{1}$ (Line 4).
-The variables $epReward$ and donedone are also initialized to track the cumulative reward and episode completion status, respectively (Lines 56).
-The while loop (Line 7) continues until the episode concludes.
-During each iteration of the while loop, an action is sampled from the policy $\pi$ (Line 8), and both the observation and action are converted into natural language representations by the captioners.
-These representations are added to the experience dataset along with the cumulative reward (Line 9).
-Subsequently, the environment is updated using the selected action, and the next observation is retrieved (Line 10).
-Finally, the cumulative reward and step index are updated accordingly (Lines 1112).
-The information stored in the experience dataset is detailed in B.
-4.4 Generating Natural Language Summaries
-We move on to addressing the second phase of Algorithm 1, covered in Lines 1517.
-To generate natural language summaries, we use the experience dataset and an LLM.
-The process begins by formatting the experience dataset into a structured format suitable for integration into a prompt.
-This format captures the sequential and contextual details of the agent’s interactions, creating a comprehensive narrative of its decision-making process (Line 15).
-The prompt design, inspired by Chain-of-Thought reasoning [47, 48], decomposes complex behavioral analysis into specific components.
-By incorporating detailed environmental parameters and quantitative evaluation guidelines, the prompt effectively guides the model’s analysis, enhancing its ability to generate insightful summaries of reinforcement learning behavior.
-This design was refined through iterative exploration, as detailed in I.
-The prompt structure consists of four main components:
-  • General Instructions: for example, “Generate a summary that captures the essence of the agent’s policy.”
-  • Environment Description: such as, “Describe the agent’s task and key aspects of the environment.”
-  • Interpretation Instructions: e.g., “Descriptions are from the agent’s perspective.”
-  • Output Instructions: for instance, “Provide a concise summary.”
-The LLM then produces the summary based on the prompt and the formatted experience dataset (Line 16).
-Algorithm 1 SySLLM
-Input: Environment $ENV$, trained policy $\pi$, captioners $C_{\text{observation}}$, $C_{\text{action}}$
-Parameter: number of episodes $N$, prompt $P$
-Output: summary $\in\mathcal{T}$
-1: Initialize experience dataset EDED
+SySLLM generating policy summaries
+Figure 2. Generating global policy summaries (Section 4.3).
+4. SySLLM Framework
+Our SySLLM (Synthesized Summary using LLMs) framework formalizes policy summarization as a two-phase process: (i) experience collection and captioning, which transforms trajectories into natural language traces stored in a Textual Experience Buffer (TEB), and (ii) abstractive summarization, which synthesizes the TEB into a concise global description of the agent’s policy using a large language model (LLM).
+Algorithms 12 specify the pipeline, while Figures 12 illustrate its two phases.
+4.1. Captioners
+The first phase converts raw trajectories into textualized experiences. For a trajectory $\tau=(o_{1},a_{1},r_{1},\ldots,o_{T})$, we define two captioning functions:
+$C_{\text{obs}}:\mathcal{O}\to\Sigma^{\star},\quad C_{\text{act}}:\mathcal{A}\to\Sigma^{\star},$
+where $\Sigma^{\star}$ denotes the set of natural language strings.
+At each step $t$, the pair $(o_{t},a_{t})$ is mapped to $(C_{\text{obs}}(o_{t}),C_{\text{act}}(a_{t}))$, yielding a textual description of the observation and the action taken.
+Observation Captioner. $C_{\text{obs}}$ produces structured descriptions of salient percepts (e.g. “The agent is facing a locked door with a key to the left.”).
+Action Captioner. $C_{\text{act}}$ verbalizes the agent’s action (e.g. “move forward,” “pick up the key.”).
+This follows prior work on language rounding jiang2019language; mirchandani2021ella.
+In practice, captioners may be rule-based, vision-language based, or hybrid; the concrete instantiation is left as a domain-specific choice and is discussed further in Section 9.
+4.2. Constructing the Textual Experience Buffer
+From the $N$ sampled episodes, we construct a Textual Experience Buffer (TEB):
+$TEB_{\pi}=\langle e_{1},e_{2},\dots,e_{L}\rangle,\quad e_{t}=\langle\hat{o}_{t},\hat{a}_{t},r_{t},epID\rangle,$
+where $\hat{o}_{t}=C_{\text{obs}}(o_{t})$ and $\hat{a}_{t}=C_{\text{act}}(a_{t})$ denote natural language captions of observations and actions.
+Each tuple records captioned observation, captioned action, reward, and episode identifier, preserving temporal coherence between steps.
+Episodes are generated by sampling from the policy $\pi$:
+$o_{1}\sim\texttt{RESET}(ENV),\quad a_{t}\sim\pi(o_{t}),\quad(o_{t+1},r_{t},done)\sim\texttt{STEP}(a_{t}),$
+repeating until $done=\texttt{True}$. At each step, both observation and action are passed through captioning functions $C_{\text{obs}}$ and $C_{\text{act}}$.
+The TEB is thus a textual analogue of a replay buffer: it aggregates multiple trajectories in a structured, language-based format that is directly consumable by LLMs, while maintaining the sequential structure necessary for policy-level reasoning.
+Further details on the fields stored in the TEB is provided in Appendix B.
+4.3. Abstractive Summarization
+Let $\mathcal{B}$ denote the space of textual experience buffers, where each $TEB_{\pi}\in\mathcal{B}$ is a finite ordered sequence of captioned experience tuples collected from executions of policy $\pi$.
+The second phase maps the TEB to a global narrative $T\in\mathcal{T}$. We define the summarizer as
+$f_{\theta}:\mathcal{B}\rightarrow\mathcal{T},\qquad T=f_{\theta}(TEB_{\pi}),$
+where $\Sigma^{\star}$ denotes the space of finite token sequences and $f_{\theta}$ is instantiated by an LLM conditioned on a structured prompt.
+Prompt Construction.
+The prompt follows a hierarchical structure, inspired by Chain-of-Thought reasoning wei2022chain; kojima2022large, which decomposes complex behavioral analysis into specific components.
+This design guides the LLM to progressively move from low-level traces in the TEB to higher-level abstractions of the policy.
+The prompt consists of:
+  1. (1)
+    General Instructions: define the summarization task.
+  2. (2)
+    Environment Context: describe task objectives and constraints.
+  3. (3)
+    Textual Experience Buffer: provide $TEB$ in structured form to preserve sequentiality.
+  4. (4)
+    Output Specification: constrain the output to a concise summary in natural language.
+Prompt design details are provided in Appendix I.
+Scalability via Hierarchical Summarization.
+A fundamental challenge in LLM-based summarization is the bounded context size: The TEB may exceed the maximum token budget $\kappa$ of the model.
+To address this, SySLLM employs a hierarchical procedure that recursively reduces the buffer until it fits within the context window.
+Formally, define a summarization operator:
+$\mathcal{S}_{\theta}:\mathcal{D}\to\Sigma^{\star},$
+where $\mathcal{D}$ is any subset of the TEB and $\Sigma^{\star}$ the space of textual summaries. If $|TEB|\leq\kappa$, we directly apply:
+$T=\mathcal{S}_{\theta}(TEB).$
+If $|TEB|>\kappa$, we partition the buffer into $M$ disjoint subsets $TEB=\{TEB^{(1)},\dots,TEB^{(M)}\}$ such that each $|TEB^{(i)}|\leq\kappa$. For each subset we compute intermediate summaries:
+$S_{i}=\mathcal{S}_{\theta}(TEB^{(i)}),\quad i=1,\dots,M,$
+and aggregate them by applying $\mathcal{S}_{\theta}$ again:
+$T=\mathcal{S}_{\theta}(\{S_{1},\dots,S_{M}\}).$
+This recursive divide-and-conquer scheme:
+$T=\textsc{HierarchicalSummarize}(TEB,\kappa)$
+ensures that SySLLM remains applicable to arbitrarily large buffers while preserving coverage across all episodes and compressing details into intermediate summaries.
+Candidate Generation and Selection.
+Once the input (original or hierarchical) fits within context, we query the LLM to generate $K$ candidate summaries $\{T_{1},T_{2},\dots,T_{K}\}$ via stochastic decoding.
+This captures variability in abstraction and phrasing.
+To select a robust final summary, each candidate is embedded into a semantic vector space using a pretrained embedding model, yielding $\phi(T_{i})\in\mathbb{R}^{d}$.
+We compute the centroid:
+$c=\tfrac{1}{K}\sum_{i=1}^{K}\phi(T_{i}),$
+and measure distances:
+$d_{i}=\lVert\phi(T_{i})-c\rVert_{2}.$
+The final summary $T^{*}$ is chosen as the median representative, i.e. the candidate closest to the median-ranked distance from the centroid:
+$T^{*}=\arg\min_{T_{i}}\;\big|\text{rank}(d_{i})-\tfrac{K}{2}\big|.$
+This selection scheme balances generality (summaries near the centroid capture broad regularities) with specificity (summaries farther away capture contextual details), yielding a consensus-style narrative.
+Algorithm 1 SySLLM Framework
+Input: Environment $ENV$, trained policy $\pi$, captioners $C_{\text{obs}}$, $C_{\text{act}}$, base prompt $P$
+Parameters: number of episodes $N$, token budget $\kappa$, #candidates $K$, embedding model $\phi(\cdot)$
+Output: Policy summary $T^{*}\in\mathcal{T}$
+1:Initialize Textual Experience Buffer $TEB\leftarrow\varnothing$
 2:for $i=1$ to NN do
-3:  Let $t\leftarrow 1$
-4:  $o_{t}\leftarrow$ ENVENV.RESET()
-5:  $epReward\leftarrow 0$
-6:  $done\leftarrow\text{False}$
-7:  while $\neg done$ do
-8:   $a_{t}\sim\pi(\cdot\mid o_{t})$
+3:  $t\leftarrow 1$; $o_{t}\leftarrow ENV.\texttt{RESET}()$; $epReward\leftarrow 0$; $done\leftarrow\textbf{False}$
+4:  while $\neg done$ do
+5:   $a_{t}\sim\pi(\cdot\mid o_{t})$
+6:   $TEB.\texttt{ADD}\big(\,C_{\text{obs}}(o_{t}),\,C_{\text{act}}(a_{t}),\,epReward,\,i\,\big)$
+7:   $(o_{t+1},r_{t+1},done)\leftarrow ENV.\texttt{STEP}(a_{t})$
+8:   $epReward\leftarrow epReward+r_{t+1}$; $t\leftarrow t+1$
+9:   $o_{t}\leftarrow o_{t+1}$
+10:  end while
+11:end for
 […]
-10:   $o_{t+1},r_{t+1},done\leftarrow$ ENVENV.STEP($a_{t}$)
-11:   $epReward\leftarrow epReward+r_{t+1}$
-12:   Let $t\leftarrow t+1$
-13:  end while
-14:end for
-15:formattedInputformattedInput $\leftarrow$ $P$ + FORMAT(EDED)
-16:summarysummary $\leftarrow$ LLM(formattedInputformattedInput)
-17:return summarysummary
-5 Implementation
-We showcase our method in five distinct environments from the MiniGrid Framework [49] as well as the challenging Crafter environment [50].
-5.1 MiniGrid
-MiniGrid is a framework for 2D based environments that feature goal-oriented tasks where agents navigate mazes, interact with objects like doors or keys, and solve challenges.
-We evaluate our approach in five distinct MiniGrid environments using seven different agents, each following a unique policy.
-5.1.1 Captioners
-To convert raw observations and actions into natural language, we designed a rule-based captioning system for the agent’s grid-based representation of the environment.
-This system systematically generates detailed descriptions of visible elements, emphasizing spatial relationships for navigation.
-Each action in the environment’s discrete action space is translated into natural language annotations, such as ‘move forward’ or ‘turn left’.
-We evaluated the effectiveness of this captioner by experimenting with and refining its design, ensuring the generated descriptions are accurate and comprehensive.
-5.1.2 Agents’ Policy Training
-We trained three qualitatively different agents for the MiniGrid-Unlock environment:
-  • Goal-directed Agent: Follows a policy designed to minimize the number of steps to unlock the door efficiently. The agent can see all grid cells within a 180-degree field of vision in front of it.
-  • Short-Sighted Agent: Follows a policy designed to minimize the number of steps to unlock the door efficiently but the agent’s sight is $3\times 3$.
-  • Random Agent: Selects actions uniformly from the action space. We chose this policy to see how the LLM summarization process handles cases where there is no clear pattern of behavior.
-Additionally, we trained four other agents in the following MiniGrid environments: Dynamic Obstacles, Lava Gap, Red-Blue Doors, and Crossing.
-Using SySLLM, we generated a policy summary for each of these agents.
-All of the MiniGrid agents are trained using the PPO algorithm [51, 52].
-An overview of agents’ performance is shown in Table 1.
-The mean reward for the first three agents in the Unlock environment shows a declining trend, which aligns with the policies that each of these agents employ.
-In addition, the success rate of the agents in the other environments remains consistently high.
-Implementation details are provided in F.
-Agent Mean Reward ± SD Mean Length Success Rate
-Unlock Goal-directed 0.73 ± 0.21 20.25 0.93
-Unlock Short-sighted 0.41 ± 0.27 44.43 0.77
-Unlock Random 0.00 ± 0.01 70.00 0.00
-Dynamic Obstacles 0.78 ± 0.06 17.20 1.00
-Lava Gap 0.90 ± 0.02 10.82 1.00
-Red Blue Doors 0.70 ± 0.26 17.06 0.88
-Crossing 0.67 ± 0.18 24.80 0.94
-Table 1: Report of agents’ performance metrics for 500 episodes for the MiniGrid environments across 3 different seeds.
-5.1.3 Generating Policy Summaries
-For each agent, we collect an experience dataset that logs the captioned observations, actions taken by the agent, and additional relevant data.
-Specifically, we log 50 episodes for each agent.
-After collecting and formatting the data, we synthesized the narratives into coherent global policy summaries for each agent using zero-shot prompting.
-We employed the gpt-4-turbo model with a temperature of $0.2$ [53].
-The prompt used is shown in C.
-In selecting the final summary, we employed a methodical approach by first generating ten distinct summaries through Line 16 of our algorithm.
-Each of these summaries was then embedded into a high-dimensional vector space using text-embedding-3-small, allowing us to quantify the semantic proximity of each summary to the others within this space.
-The centroid of these embeddings was calculated to identify the geometric mean of the summaries, representing the central tendency of the semantic dimensions explored by the model.
-We chose the summary that is the median (5th) closest to the centroid.
-This position represents a solution to the trade-off between generality and specificity.
-Summaries closer to the centroid tend to include more general statements that apply broadly across different scenarios, while those further away incorporate more specific details unique to particular contexts.
-An example of a selected summary is provided in D.
-5.2 Crafter
-Secondly, we evaluate SySLLM in the Crafter environment, a 2D adaptation of Minecraft.
-Similar to Minecraft, Crafter features a procedurally generated, partially observable world where agents can gather resources and craft artifacts structured within an achievement tree.
-This tree outlines all possible achievements along with their prerequisites.
-Crafter lacks a singular main objective, allowing agent progress to be measured by the achievements it unlocks.
-5.2.1 Captioners
-To convert the environment observation representations and actions in Crafter into natural-language format, we implemented a captioning system tailored to the game’s mechanics.
-The Observation Captioner component generates comprehensive descriptions of the player’s current state, including inventory contents, achievements, and spatial relationships to nearby resources and threats.
-For example, a typical observation might be transformed from raw data into ”The player is facing right.
-Inventory: 2 wood, 1 stone.
-A zombie is ahead…”.
-The spatial awareness system maps objects to their relative coordinates.
-The Action Captioner component complements this by translating action indices into meaningful descriptions like “place_table”, “move_right” or “make_wood_sword”, providing context for the agent’s decisions.
-Figure 3 provides an example of how the captioners translate the agent’s trajectory into natural language descriptions.
-Figure 3: This figure illustrates four steps from a trajectory of the Resource-Collector agent in the Crafter environment, alongside their corresponding natural language captions generated using the observation and action captioners, $C_{\text{observation}}$ and $C_{\text{action}}$.
-For each step, the captions describes the agent’s inventory status, the object currently in front of it, and the next action selected by the agent.
+13:return $T$
+Algorithm 2 HierarchicalSummarize
+Input: textual experience buffer subset $D$, token budget $\kappa$, prompt $P$, #candidates $K$, embedding model $\phi$
+Output: summary $T\in\mathcal{T}$
+1:if $\texttt{TOKENS}(\texttt{FORMAT}(D))\leq\kappa$ then
+2:  $X\leftarrow P\;+\;\texttt{FORMAT}(D)$
+3:  $\{T_{1},\ldots,T_{K}\}\leftarrow\texttt{LLM.SAMPLE}(X,K)$
+4:  $E_{i}\leftarrow\phi(T_{i})\in\mathbb{R}^{d}\quad\forall i\in\{1,\ldots,K\}$
+5:  $c\leftarrow\frac{1}{K}\sum_{i=1}^{K}E_{i}$
+6:  $d_{i}\leftarrow\lVert E_{i}-c\rVert_{2}\quad\forall i$;   $j^{\star}\leftarrow\arg\min_{j}\left|\,\texttt{rank}(d_{j})-\frac{K}{2}\,\right|$
+7:  return $T_{j^{\star}}$
+8:else
+9:  $\{D^{(1)},\ldots,D^{(M)}\}\leftarrow\texttt{PARTITION}(D,\kappa)$ $\triangleright$ disjoint, each fits
+10:  for $m=1$ to $M$ do
+11:   $S_{m}\leftarrow\textsc{HierarchicalSummarize}(D^{(m)},\kappa,P,K,\phi)$
+12:  end for
+[…]
+14:end if
+5. Experimental Setup
+We evaluate SySLLM across controlled reinforcement learning environments that capture both simple and complex agent behaviors.
+Specifically, we apply our framework to five environments from the MiniGrid suite chevalier2021minimalistic and the more challenging Crafter environment hafner2021crafter.
+These environments were selected to span a range of task structures, observation modalities, and policy complexities.
+We use the gpt-4-turbo model with a temperature of 0.5 achiam2023gpt, and the text-embedding-3-small model openai2024embedding as the embedding function.
+5.1. MiniGrid
+MiniGrid is a grid-world framework where agents perform goal-directed navigation and object-interaction tasks under partial observability.
+We instantiate seven agents across five environments, ensuring diversity in both policy performance and behavioral style.
+Captioners.
+To construct the Textual Experience Buffer (TEB), we implement a rule-based captioning system that maps raw grid observations and discrete actions into structured natural-language descriptions.
+Observation Captioner ($C_{\text{obs}}$): generates textual descriptions of visible elements, including object types (e.g., keys, doors, obstacles) and their spatial relationships relative to the agent.
+Action Captioner ($C_{\text{act}}$): translates the discrete action set (e.g., turn_left, move_forward) into natural-language strings.
+Agents and Training We trained three agents with qualitatively distinct policies in the MiniGrid-Unlock environment:
+  • Goal-directed agent: wide $7\times 7$ observation window, optimized to minimize the steps to unlock the door.
+  • Short-sighted agent: restricted $3\times 3$ observation window, leading to more myopic strategies.
+  • Random agent: selects actions uniformly at random, providing a non-structured behavioral baseline.
+In addition, we trained one agent each in four further environments: Dynamic Obstacles, Lava Gap, Red-Blue Doors, and Crossing.
+All MiniGrid agents were trained using PPO schulman2017proximal; stable-baselines3 for 1M timesteps per seed for three random seeds.
+Agent performance statistics are reported in Table 1.
+The complete hyperparameters are provided in the Appendix F.
+TEB Collection and Summarization.
+For each agent, we collect $50$ evaluation episodes to construct the TEB.
+Each buffer is formatted into the structured prompt and processed as described in Section 4.3, where candidate summaries are generated via stochastic decoding and post-processed for abstractive summarization.
+Prompt templates used for the MiniGrid suite domain are listed in Appendix C.
+5.2. Crafter
+Crafter is a 2D, partially observable world inspired by Minecraft, featuring procedurally generated maps, resource gathering, crafting, and an achievement tree that defines agent progress.
+Captioners To adapt the captioning system to Crafter, we extend the MiniGrid captioner to encode: Observation Captioner: inventory contents, spatial relations to nearby resources and threats, current health and stamina, and unlocked achievements.
+Action Captioner: maps the Crafter action set into textual forms, e.g., “move_right”, “place_table”, “make_wood_sword”.
+Figure 3 shows an example trajectory and its textualized representation.
+Agents and Training We train two agents with distinct behavior:
+  • Resource-Collector agent: trained with DreamerV3 hafner2023dreamerv3, capable of sustained survival, resource collection, and the crafting of basic tools.
+  • Random agent: uniformly samples from the action set, serving as a baseline with no structured policy.
+TEB Collection and Summarization.
+For each agent, we log $5$ evaluation episodes.
+Due to the long horizon of Crafter, the TEB for each episode can exceed the LLM’s context length.
+In this case, we apply the hierarchical summarization strategy described in Section 4.3.
+Prompt templates used for the Crafter domain are listed in Appendix E.
+[…]
+Crafter captioning
+Figure 3.
+Four steps from a trajectory of the Resource-Collector agent in the Crafter environment, alongside their corresponding captions generated using the observation and action captioners.
+For each step, the captions describe the agent’s inventory status, the object currently in front of it, and the next action selected by the agent.
 A textual representation of the visible grid (highlighted in blue) is also included to reflect the agent’s local perception.
 Additionally, all unique achievements unlocked by the agent throughout the trajectory are summarized in red.
-5.2.2 Agents’ Policy Training
-We trained two qualitatively different agents for the Crafter environment.
-The first, named Resource-collector Agent, is a well-trained agent capable of navigating in Crafter with efficiency and skill.
-The agent is capable to search for food and water, defend against monsters, collect materials, and craft tools while unlocking semantically meaningful achievements along the achievement tree.
-This agent was trained using the DreamerV3 algorithm [54].
-The second agent, a Random Agent, selects actions uniformly at random from the available action space without any learned strategy.
-Implementation details are provided in F.
-5.2.3 Generating Policy Summaries
-Similar to the MiniGrid environments, we gather an experience dataset for each agent, recording captioned observations, the actions taken by the agent, and other relevant information.
-Specifically, we log five episodes per agent.
-Due to context window limitations, each episode is summarized individually, and the resulting episode summaries are then combined into a single comprehensive summary.
-The prompts used are shown in E.
-5.3 Insights From Summaries
-Using the SySLLM method, we produced summaries that highlight essential aspects of each agent’s decision-making and behavior.
-This section outlines key insights derived from these summaries, offering a clear depiction of the agents’ policies and significant action patterns.
-Table 3 presents an overview of these insights for the MiniGrid agents, while Table 4 provides an overview for the Crafter agents.
-These insights showcase the alignment of SySLLM-generated summaries with task objectives and illustrate how agents perform across diverse environments.
-For the goal-directed agent in the Unlock environment, Insight (2) states that the agent “demonstrates a pattern of turning towards the nearest key or door once identified, suggesting a straightforward heuristic of minimizing distance to the target.” This reflects the agent’s efficient and consistent behavior, dynamically adjusting its path to achieve objectives with minimal steps.
-Insight (3) quantifies this efficiency, noting that episodes are completed within 15-25 steps on average, often achieving near-maximum rewards.
-These observations are validated by performance metrics in Table 1, which report a mean episode length of 20.25 and a high mean reward of 0.71.
-Together, these results highlight SySLLM’s ability to capture both qualitative patterns and quantitative performance metrics.
-In the Dynamic Obstacles environment, SySLLM captures the agent’s ability to navigate challenges effectively.
-Insight (1) states that the agent “consistently demonstrates the ability to avoid obstacles (blue balls) by making turns or moving forward when the path is clear.” This highlights its proficiency in avoiding collisions, a critical success factor.
-However, Insight (3) notes inconsistent behavior when navigating closely spaced obstacles, with the agent occasionally taking longer routes or making unnecessary turns.
-These observations demonstrate SySLLM’s ability to provide a nuanced understanding of both the agent’s strengths in straightforward scenarios and its struggles in more complex configurations.
-For the Resource-collector agent in the Crafter environment, SySLLM effectively captures the agent’s behavior as a resource-collector, emphasizing its focus on gathering essential materials.
-Insight (1) highlights that “The agent consistently prioritizes acquiring fundamental resources such as wood, stone, and drink, which are critical for crafting tools and sustaining basic survival metrics.” Furthermore, SySLLM captures the agent’s achievement-oriented tendencies while identifying limitations in unlocking advanced achievements.
-As stated in Insight (4), “The agent reliably achieves milestones related to resource collection and basic tool crafting but demonstrates difficulty in attaining more advanced achievements, suggesting potential areas for improvement.” These observations are validated by performance metrics presented in Table 2, which report a mean reward of 10.42—indicating relatively high performance—and a mean of 11.33 unique achievements unlocked, reinforcing the agent’s proficiency in foundational tasks while highlighting room for growth in advanced capabilities.
-Agent Mean Reward ± SD Mean Length Mean Unique Achievements
-Resource-Collector 10.43 ± 2.11 234.6 11.33
-Random 1.39 ± 1.19 164.4 2.29
-Table 2: Report of agents’ performance metrics for 100 episodes for the Crafter environment across 3 different seeds.
-Agent Insights from Summaries
-Unlock
-Goal–
-directed
-(1) The agent effectively identifies keys and adjusts its path based on their relative position, shifting focus to unlocking the door.
-(2) It consistently turns towards the nearest key or door, minimizing distance, which remains consistent across episodes.
-(3) The agent completes episodes efficiently, averaging 15–25 steps with near-maximal cumulative rewards.
-Unlock
-Short–
-sighted
-(1) The agent follows a right-wall method, moving forward until encountering an obstacle before turning.
-(2) It identifies keys and doors efficiently, maneuvering toward and using them correctly.
-(3) Decisions are heavily influenced by its immediate field of vision, reacting only to nearby objects.
-Unlock
-Random
-(1) The agent exhibits unstructured behavior, often repeating unnecessary actions.
-(2) It frequently toggles doors multiple times or picks up and drops keys without using them effectively.
-Lava Gap (1) The agent consistently avoids lava, demonstrating awareness of environmental hazards.
-(2) Upon encountering an obstacle, it either turns or moves in the opposite direction.
+Env. Agent Mean Reward ± SD Mean Length Success / Achievements
+MiniGrid Unlock Goal-directed 0.73 ± 0.21 20.25 Success 0.93
+Unlock Short-sighted 0.41 ± 0.27 44.43 Success 0.77
+Unlock Random 0.00 ± 0.01 70.00 Success 0.00
+Dynamic Obstacles 0.78 ± 0.06 17.20 Success 1.00
+Lava Gap 0.90 ± 0.02 10.82 Success 1.00
+Red Blue Doors 0.70 ± 0.26 17.06 Success 0.88
+Crossing 0.67 ± 0.18 24.80 Success 0.94
+Crafter Resource-Collector 10.43 ± 2.11 234.6 11.33 achievements
+Random 1.39 ± 1.19 164.4 2.29 achievements
+Table 1.
+Performance metrics across MiniGrid (500 episodes, 3 seeds) and Crafter (100 episodes, 3 seeds).
+For MiniGrid, performance is measured by success rate; for Crafter, by the number of unique achievements unlocked.
+6. Illustrative Policy Summaries
+To illustrate the summaries produced by SySLLM, we present structured case studies that connect textual motifs to quantitative evidence and then expand them into detailed narrative accounts.
+This combination shows that the framework produces explanations that are not only linguistically coherent but also grounded in the actual behavior of the agents.
+A full summary of the Unlock Goal-Directed agent is included in Appendix D.
+Structured Case Studies.
+Table 2 reports representative agents in MiniGrid and Crafter.
+For each, we show: (i) the central claim extracted by SySLLM, and (ii) the quantitative metrics that support or refute this claim.
+The alignment between narrative motifs and behavioral statistics illustrates that SySLLM captures recurring strategies and limitations in a manner consistent with ground-truth agent performance.
+For example, the Unlock Goal-Directed agent’s motif of “turning toward the nearest key or door” is supported by a $0.93$ success rate and shorter episode lengths, while the Random agent’s “lack of coherent strategy” corresponds to maximal episode lengths and zero success rate.
+In Crafter, the Resource-Collector agent’s emphasis on survival resources is matched by high reward and achievement counts, whereas the Random agent’s incoherence is validated by minimal achievement progression.
+Env. Agent SySLLM Summary Claim Quantitative Alignment
+MiniGrid Unlock Goal-directed Turns toward nearest key/door Success $0.93$, mean length $20.25$
+Unlock Random No clear strategy Success $0.00$, max length episodes
+Dynamic Obstacles Avoids obstacles, occasional inefficiency Success $1.00$, higher variance in length
+Lava Gap Jumps gap consistently Success $1.00$, mean length $10.82$
+Crafter Resource-Collector Focused, survival-oriented strategy; Unlocks early achievements Reward $10.43$, $11.33$ unique achievements
+Random Sporadic, incoherent progression Reward $1.39$, $2.29$ unique achievements
+Table 2.
+Structured case studies of SySLLM summaries across environments.
+Extracted motifs are validated against quantitative metrics, showing alignment between narrative claims and behavioral statistics.
+Expanded Narrative Insights.
+Beyond single-sentence claims, SySLLM produces multi-faceted descriptions of agent behavior.
+Figure 4 presents detailed motifs across environments, highlighting how the framework abstracts local decisions into global patterns.
+These narratives capture not only the strengths of agents (e.g.
+consistent lava avoidance, structured door-order strategies) but also nuanced inefficiencies (e.g., unnecessary turns in dense obstacle scenarios, repetitive failed crafting attempts in Crafter).
+Such fine-grained insights demonstrate SySLLM’s ability to diagnose effective heuristics and characteristic failure modes.
+Unlock – Goal-directed
+  1. (1)
+    The agent effectively identifies keys and adjusts its path based on their relative position, shifting focus to unlocking the door.
+  2. (2)
+    It consistently turns towards the nearest key or door, minimizing distance, which remains consistent across episodes.
+  3. (3)
+    The agent completes episodes efficiently, averaging 15–25 steps with near-maximal cumulative rewards.
+Unlock – Short-sighted
+  1. (1)
+    The agent follows a right-wall method, moving forward until encountering an obstacle before turning.
+  2. (2)
+    It identifies keys and doors efficiently, maneuvering toward and using them correctly.
+  3. (3)
+    Decisions are heavily influenced by its immediate field of vision, reacting only to nearby objects.
+Unlock – Random
+  1. (1)
+    The agent exhibits unstructured behavior, often repeating unnecessary actions.
+  2. (2)
+    It frequently toggles doors multiple times or picks up and drops keys without using them effectively.
+Lava Gap
+  1. (1)
+    The agent consistently avoids lava, demonstrating awareness of environmental hazards.
+  2. (2)
+    Upon encountering an obstacle, it either turns or moves in the opposite direction.
 Red–Blue
-Doors
-(1) The agent prioritizes opening the red door before the blue door, optimizing reward accumulation.
-(2) It successfully interacts with doors in a structured sequence, adhering to task constraints.
-Crossing (1) The agent moves towards the green goal once it enters its field of vision, adjusting its path accordingly.
-(2) It avoids collisions with walls through timely directional changes.
-Dynamic
-Obstacles
-(1) The agent effectively avoids moving obstacles (blue balls) by adjusting its movement.
-(2) It identifies objects in its field of vision and makes informed navigation decisions.
-(3) In dense obstacle scenarios, occasional inefficiencies or unnecessary turns are observed.
-Table 3: Insights from agents’ SySLLM summaries in the MiniGrid environments.
-Agent Insights from Summaries
-Crafter
-Resource-collector
-(1) Strong Focus on Resource Collection: The agent consistently prioritizes gathering essential resources such as wood, stone, and drink, which are foundational for crafting tools and maintaining basic survival metrics.
-(2) Effective Basic Tool Crafting: Regular crafting of basic tools like wood pickaxes and swords enables the agent to enhance resource collection and engage in occasional combat.
-(3) Achievement Unlocking: The agent reliably unlocks achievements related to resource collection and basic tool crafting but struggles with more advanced achievements, highlighting a potential area for improvement.
-(4) Moderate Combat Engagement: The agent occasionally engages with zombies, using crafted tools for defense, showing moderate adaptability to threats but limited combat readiness overall.
-(5) Predictable Behavior: Episodes are characterized by high consistency in resource collection and basic tool crafting actions, with low variance across episodes.
-(6) Inconsistent Health and Exploration Management: While the agent effectively manages drink levels, it shows less consistency in food and health management and sacrifices exploration efficiency for achievement unlocking.
-Crafter
-Random
-(1) Ineffective Crafting and Resource Management: The agent frequently attempts crafting without the necessary resources or understanding of prerequisites, leading to repeated failures and minimal progress in achieving complex objectives.
-(2) Poor Survival Strategy: The agent consistently demonstrates ineffective survival behavior, including health depletion and poor management of food and drink levels, which hampers its ability to sustain itself in the game environment.
-(3) Limited Achievement Progression: While the agent reliably unlocks basic achievements like ‘wake_up’ and ‘collect_sapling’, it struggles to achieve more complex milestones that require crafting, resource management, or combat engagement.
-(4) Repetitive and Ineffective Actions: Episodes are marked by high frequencies of ‘noop’ actions and repetitive failed attempts at crafting, reflecting a lack of strategic adaptation and learning from past failures.
-(5) Lack of Combat Engagement: Lack of Combat Engagement: The agent shows minimal engagement with combat mechanics and fails to defend effectively against threats such as zombies and skeletons.
-(6) Predictable Behavior: Across episodes, the agent exhibits consistent, ineffective patterns of action, suggesting significant limitations in its decision-making processes and adaptability.
-Table 4: Insights from agents’ SySLLM summaries in the Crafter environment.
-6 Expert Evaluation of Summaries
-To evaluate the quality and alignment of SySLLM summaries, we conducted evaluations with RL experts.
-Specifically, six graduate students actively engaged in RL research with experience in training agents were recruited.
-These experts were divided into two groups: Experts 1, 2, and 3 evaluated the MiniGrid-Unlock goal-directed and short-sighted agents, while Experts 4, 5, and 6 evaluated the the rest of the MiniGrid agents.
-Experts 1,2, and 6 evaluated the agents from the Crafter environment.
-Each expert was shown videos approximately 120 seconds long, depicting the agents’ behaviors in the environment.
-Based on these videos, the experts were tasked with summarizing the agents’ behaviors.
-To ensure a fair comparison with SySLLM-generated summaries, we provided the experts with instructions that aligned closely with the prompts used for the language model (see G for detailed instructions).
-To evaluate how much of the experts’ observations were captured by the SySLLM summaries, we used a recall metric.
-This metric quantified the extent to which the SySLLM summary covered the points identified by the experts.
-Each expert summary and the SySLLM summary were broken down into lists of key points, and a scoring system was applied: 1 point for a match with a SySLLM key point, 0.5 points for a partial match, and 0 points for no match.
-The recall scores, as shown in Table 5, ranges from 0.687 for the Unlock goal-directed agent, indicating moderate agreement, to 0.914 for the Crossing agent, reflecting a near-complete capture of expert insights.
-The overall mean recall score across all agents is 0.840, demonstrating substantial coverage of the experts’ points by the SySLLM summaries.
-In the second evaluation phase, the experts were presented with key points from the SySLLM summaries that they had not mentioned in their own summaries.
-They were tasked with classifying these points as Matched, Partially Matched, or Not Matched based on their agreement with the content.
-Using the same scoring system as for recall (Matched = 1, Partially Matched = 0.5, Not Matched = 0), we calculated a precision score for each summary.
-The precision score measures how many of the SySLLM summary points were validated as correct by the experts.
-Importantly, the precision calculation includes all SySLLM key points, encompassing both the points classified in this phase and those already evaluated during the recall assessment.
-The precision scores, as shown in Table 5, range from 0.769 (Dynamic Obstacles agent) to 0.864 (Short-Sighted agent), with an overall mean precision score of 0.839 across all agents.
-In addition to quantifying correctness, this metric offers a way to evaluate potential hallucinations by identifying points that may not align with expert observations.
-The high precision scores suggest that hallucinations were minimal in our evaluations.
-While instances of hallucinated content in SySLLM summaries were rare, they underscore the inherent challenges in summarizing RL agent behaviors using LLMs.
-One example from the Crossing agent summary: “The agent frequently checks for walls in its path and adjacent tiles.”.
-This point is classified as “not matched” by all experts, indicating a possible hallucination.
-To assess the reliability of expert annotations, we report both percentage agreement and Gwet’s AC1–metrics appropriate for imbalanced labeling tasks.
-The mean agreement reached 70%, and Gwet’s AC1 indicated substantial reliability (mean = 0.72), reflecting strong consistency across annotators.
-These results not only validate the quality of the annotations, but also provide important context for interpreting model performance: if human experts do not fully agree, perfect alignment between model and human labels is neither expected nor realistic.
-In this light, the model’s ability to approximate expert consensus is particularly notable.
-Taken together, the integration of recall and precision metrics demonstrates SySLLM’s capability to produce comprehensive and accurate summaries that closely align with expert observations.
-Agent Expert Recall Precision Mean Recall Mean Precision
-Unlock
-Goal-Directed
-1 0.500 0.864 0.687 0.864
-2 0.643 0.864
-3 0.917 0.864
-Unlock
-Short-Sighted
-1 0.800 0.846 0.878 0.839
-2 0.833 0.807
-3 1.000 0.923
-Dynamic
-Obstacles
-4 0.583 0.692 0.739 0.769
-5 0.833 0.692
-6 0.800 0.923
-Lava Gap
-4 0.667 0.769 0.794 0.811
-5 0.786 0.846
-6 0.929 0.818
-Red Blue
-Doors
-4 0.857 0.767 0.871 0.834
-5 0.857 0.867
-6 0.900 0.867
+  1. (1)
+    The agent prioritizes opening the red door before the blue door, optimizing reward accumulation.
+  2. (2)
+    It successfully interacts with doors in a structured sequence, adhering to task constraints.
 Crossing
-4 0.750 0.731 0.914 0.795
-5 0.917 0.808
-6 1.000 0.846
-Crafter
-Resource-Collector
-7 0.938 0.893 0.931 0.871
-8 0.938 0.857
-9 0.917 0.864
-Crafter
-Random
-7 0.929 0.917 0.902 0.929
-8 1.000 0.958
-9 0.778 0.857
-Table 5: Recall and Precision of Expert Insights for SySLLM.
-7 User Study
-To evaluate the usefulness of the SySLLM summaries, we conducted a user study comparing SySLLM summaries to HIGHLIGHTS-DIV (hereafter referred to as “HIGHLIGHTS”) summaries, which is a standard video-based benchmark in XRL [17].
-For the study, we focused on the three agents from the MiniGrid Unlock environment: goal-directed agent, short-sighted agent, and random agent.
-These agents were chosen to represent qualitatively distinct behaviors, ensuring a diverse set of policies for participants to evaluate.
-This variety allowed us to better assess how well the summaries captured different decision-making strategies and provided interpretable insights into the agents’ behavior.
-For completeness, we provide an intuitive explanation of HIGHLIGHTS in A.
-We assessed participants’ subjective preferences as well as their ability to identify agent behavior based on the summaries.
-7.1 Procedure
-Participants were first shown a tutorial explaining the rules of the Unlock environment.
-They then had to pass a quiz ensuring they read and understood the rules.
-Next, they were asked to complete two different tasks.
-Task 1: Summary Preferences
-In the first task, participants evaluated and compared summaries of agents based on their policies.
-Initially, they watched a 120-second video showcasing full episodes to familiarize themselves with the agent’s policy.
-Following this, participants were presented with the first summary, either from the SySLLM or HIGHLIGHTS condition (see Section 7.2), and rated their agreement on a series of metrics using a 7-point Likert scale.
-The metrics, adapted from Hoffman’s explanation satisfaction questionnaire [55], included Clarity, Understandability, Completeness, Satisfaction, Usefulness, Accuracy, Improvement, and Preference.
-These adaptations were tailored to the study’s specific needs and excluded irrelevant items (see G for the full scale).
-Participants then evaluated the second summary, corresponding to the alternative condition (HIGHLIGHTS or SySLLM), using the same set of questions.
-This allowed for a direct comparison between the two summary types.
-After assessing both summaries, participants answered additional questions that compared the summaries and expressed their preferences, highlighting which summary they found more helpful and explaining their reasons.
-Task 2: Agent Identification
-While the first task evaluated participants’ subjective perceptions of the summaries, the second task assessed their objective ability to match a given summary to the correct agent.
-Participants were presented with summaries, each linked to either the SySLLM or HIGHLIGHTS condition, and shown three short experience videos (approximately 20 seconds each) depicting different agents’ actions in the environment.
-They were asked to identify which video corresponded to the behavior described in the summary.
-Participants indicated their chosen video and rated their confidence on a 1-7 Likert scale, along with justifications for their decision in an open text format.
-To keep the experiment concise and focused on key comparisons, we set up three matching questions—labeled $Q1$, $Q2$, and $Q3$.
-Each question featured a summary of a different agent: the goal-directed agent in $Q1$, the random agent in $Q2$, and the short-sighted agent in $Q3$.
-Depending on the condition to which a participant was assigned, they encountered either a text-based summary (SySLLM) or a video summary (HIGHLIGHTS) for these agents.
-In each question, alongside the video of the associated agent, participants were also shown two additional videos featuring different, unrelated agents.
-The task for the participants was to correctly identify the agent that matched the given summary.
-Success was measured by the participant’s ability to accurately associate the correct agent with its respective summary.
-7.2 Experimental Conditions
-Participants were divided into four conditions, varying by summary type (SySLLM or HIGHLIGHTS) and agent type (goal-directed or short-sighted).
-For Task 1, we used a within-subject design where participants viewed summaries of an agent’s behavior in two sequences.
-To counter ordering effects, one group received the SySLLM summary first and the HIGHLIGHTS summary second, while the other group received the reverse order.
-In Task 2, we employed a between-subject design.
-Participants viewed only SySLLM or HIGHLIGHTS summaries, corresponding to the sequence they were exposed to in Task 1.
-This setup ensured that participants experienced one of four unique combinations of summary and agent types in a controlled manner.
-Details of these conditions are shown in Table 6.
-Benchmark
-For generating the HIGHLIGHTS summaries, we used the following parameters: 300 traces (episodes), a context length of 5, and a total of 20 highlights to effectively capture key behaviors and the necessary context within the Unlock environment.
-These choices ensure that the summaries provide a detailed yet concise depiction of critical actions, appropriately matching the environment’s scale and complexity.
+  1. (1)
+    The agent moves towards the green goal once it enters its field of vision, adjusting its path accordingly.
+  2. (2)
+    It avoids collisions with walls through timely directional changes.
+Dynamic Obstacles
+  1. (1)
+    The agent effectively avoids moving obstacles (blue balls) by adjusting its movement.
+  2. (2)
+    It identifies objects in its field of vision and makes informed navigation decisions.
+  3. (3)
+    In dense obstacle scenarios, occasional inefficiencies or unnecessary turns are observed.
+Crafter – Resource-collector
+  1. (1)
+    Strong Focus on Resource Collection: The agent consistently prioritizes gathering essential resources such as wood, stone, and drink, which are foundational for crafting tools and maintaining basic survival metrics.
+  2. (2)
+    Effective Basic Tool Crafting: Regular crafting of basic tools like wood pickaxes and swords enables the agent to enhance resource collection and engage in occasional combat.
+  3. (3)
+    Achievement Unlocking: The agent reliably unlocks achievements related to resource collection and basic tool crafting but struggles with more advanced achievements, highlighting a potential area for improvement.
+  4. (4)
+    Moderate Combat Engagement: The agent occasionally engages with zombies, using crafted tools for defense, showing moderate adaptability to threats but limited combat readiness overall.
+  5. (5)
+    Predictable Behavior: Episodes are characterized by high consistency in resource collection and basic tool crafting actions, with low variance across episodes.
+  6. (6)
+    Inconsistent Health and Exploration Management: While the agent effectively manages drink levels, it shows less consistency in food and health management and sacrifices exploration efficiency for achievement unlocking.
+Crafter – Random
+  1. (1)
+    Ineffective Crafting and Resource Management: The agent frequently attempts crafting without the necessary resources or understanding of prerequisites, leading to repeated failures and minimal progress in achieving complex objectives.
+  2. (2)
+    Poor Survival Strategy: The agent consistently demonstrates ineffective survival behavior, including health depletion and poor management of food and drink levels, which hampers its ability to sustain itself in the game environment.
+  3. (3)
+    Limited Achievement Progression: While the agent reliably unlocks basic achievements like wake_up and collect_sapling, it struggles to achieve more complex milestones that require crafting, resource management, or combat engagement.
+  4. (4)
+    Repetitive and Ineffective Actions: Episodes are marked by high frequencies of ‘noop‘ actions and repetitive failed attempts at crafting, reflecting a lack of strategic adaptation and learning from past failures.
+  5. (5)
+    Lack of Combat Engagement: The agent shows minimal engagement with combat mechanics and fails to defend effectively against threats such as zombies and skeletons.
+  6. (6)
+    Predictable Behavior: Across episodes, the agent exhibits consistent, ineffective patterns of action, suggesting significant limitations in its decision-making processes and adaptability.
+[…]
+Insights from agents’ SySLLM summaries in the MiniGrid environments.
+Figure 4. Insights from agents’ SySLLM summaries in the MiniGrid environments.
+Together, these examples demonstrate that abstractive textual summaries can function as faithful and interpretable accounts of reinforcement learning policies.
+7. Expert Evaluation of Summaries
+We complement the quantitative performance analysis with an expert-based evaluation of SySLLM summaries.
+The goal is to assess how well the generated summaries capture the behavioral motifs observed by human experts and to quantify correctness while identifying potential hallucinations.
+Evaluation Protocol.
+We recruited six graduate students with research experience in the training and evaluation of RL agents.
+The experts were divided into two groups: Experts 1–3 annotated the MiniGrid-Unlock (goal-directed and short-sighted) agents, while Experts 4–6 annotated the remaining MiniGrid agents.
+In addition, Experts 1, 2, and 6 annotated the Crafter agents.
+Each expert was shown a 120-second video per agent, depicting representative trajectories.
+Based on these trajectories, the experts were instructed to produce textual summaries using the same SySLLM prompting guidelines (see Appendix G).
+This alignment ensures comparability between expert- and model-generated summaries.
+Scoring Framework.
+To compare SySLLM summaries $S_{\text{LLM}}$ with expert summaries $\{S_{\text{exp}}^{j}\}_{j=1}^{m}$, we decompose both into sets of atomic key points, denoted $\mathcal{K}_{\text{LLM}}$ and $\mathcal{K}_{\text{exp}}^{j}$, respectively.
+Semantic equivalence is assessed at the level of atomic propositions, where each key point expresses a single behavioral claim.
+Following standard practice in evaluating human-interpretable explanations (hoffman2018metrics), propositions are manually extracted and matched without embedding-based or heuristic similarity measures.
+Each SySLLM key point is labeled as a full match, partial match, or unsupported relative to the expert set.
+This procedure yields a transparent and reproducible matching process, with substantial inter-annotator agreement (AC1 = 0.72).
+The full annotation protocol is provided in the appendix.
+Each pairwise comparison is scored as:
+$\text{match}(k_{\text{LLM}},k_{\text{exp}})=\begin{cases}1&\text{if semantically equivalent},\\ 0.5&\text{if partially overlapping},\\ 0&\text{otherwise}.\end{cases}$
+Recall. Recall measures the extent to which the model summary covers expert-identified key points:
+$\text{Recall}=\frac{1}{m}\sum_{j=1}^{m}\frac{\sum_{k\in\mathcal{K}_{\text{exp}}^{j}}\max_{k^{\prime}\in\mathcal{K}_{\text{LLM}}}\text{match}(k,k^{\prime})}{\lvert\mathcal{K}_{\text{exp}}^{j}\rvert}$.
+Precision.
+Precision measures the correctness of SySLLM key points relative to expert judgments.
+For each expert $j$, annotators are shown the set difference $\mathcal{K}_{\text{LLM}}\setminus\mathcal{K}_{\text{exp}}^{j}$ and asked to label each key point as Matched, Partially Matched, or Not Matched.
+Precision is defined as:
+$\text{Precision}=\frac{\sum_{k\in\mathcal{K}_{\text{LLM}}}\max_{j}\text{match}(k,\mathcal{K}_{\text{exp}}^{j})}{\lvert\mathcal{K}_{\text{LLM}}\rvert}$.
+This formulation captures both coverage (recall) and correctness (precision), while enabling explicit identification of hallucinated content through unmatched model key points.
+Results.
+Table 3 reports per-agent recall and precision.
+Recall scores range from 0.687 (Unlock goal-directed) to 0.914 (Crossing), with a mean of 0.840, demonstrating substantial overlap with expert-identified points.
+Precision scores range from 0.769 (Dynamic Obstacles) to 0.864 (Short-Sighted), with a mean of 0.839, indicating minimal hallucination.
+For example, in the Crossing environment, the point “The agent frequently checks for walls in its path and adjacent tiles” was rejected by all experts, illustrating a rare hallucination.
+Inter-Annotator Agreement.
+To ensure the reliability of expert annotations, we calculated both raw agreement (percentage of identical match scores across experts) and Gwet’s AC1 coefficient, which is robust to class imbalance in categorical judgments.
+Across all agents, the mean raw agreement was 70%, while the mean AC1 reached 0.72, indicating substantial inter-rater reliability.
+This ensures that the observed recall/precision metrics are not artifacts of inconsistent annotations.
+Taken together, these results show that SySLLM summaries exhibit both high coverage and correctness relative to expert annotations, with recall and precision consistently above 0.8.
+Agent Expert Recall Precision Mean
+Unlock Goal-Directed E1 0.500 0.864 R = 0.687, P = 0.864
+E2 0.643 0.864
+E3 0.917 0.864
+Unlock Short-Sighted E1 0.800 0.846 R = 0.878, P = 0.839
+E2 0.833 0.807
+E3 1.000 0.923
+Dynamic Obstacles E4 0.583 0.692 R = 0.739, P = 0.769
+E5 0.833 0.692
+E6 0.800 0.923
+Lava Gap E4 0.667 0.769 R = 0.794, P = 0.811
+E5 0.786 0.846
+E6 0.929 0.818
+Red-Blue Doors E4 0.857 0.767 R = 0.871, P = 0.834
+E5 0.857 0.867
+E6 0.900 0.867
+Crossing E4 0.750 0.731 R = 0.914, P = 0.795
+E5 0.917 0.808
+E6 1.000 0.846
+Crafter Resource-Collector E7 0.938 0.893 R = 0.931, P = 0.871
+E8 0.938 0.857
+E9 0.917 0.864
+Crafter Random E7 0.929 0.917 R = 0.902, P = 0.929
+E8 1.000 0.958
+E9 0.778 0.857
+Table 3. Per-expert recall and precision scores for SySLLM summaries, with aggregated per-agent means. $R$ = Recall, $P$ = Precision. Overall averages across all agents: $R=0.840$, $P=0.839$.
+8. User Study
+We conducted a controlled user study to evaluate the usefulness of SySLLM summaries compared to HIGHLIGHTS-DIV (HIGHLIGHTS) amir2018highlights, a standard demonstration-based benchmark in XRL which selects a set of high importance and diverse execution trajectories.
+The study assessed both subjective preferences and objective task performance.
+We focus on three qualitatively distinct agents from the MiniGrid Unlock environment (goal-directed, short-sighted, and random), ensuring the diversity of policies from structured strategies to noisy behaviors.
+For completeness, we provide a short description of HIGHLIGHTS in Appendix A.
+Experimental Design.
+The study used a mixed design with two tasks.
+Task 1 (Preferences) followed a within-subject setup: each participant evaluated both modalities (SySLLM and HIGHLIGHTS), with order counterbalanced to mitigate ordering effects.
+Task 2 (Identification) followed a between-subject setup: Participants viewed only one modality, aligned with their Task 1 order.
+This produced four experimental conditions that varied by summary modality and agent type, as shown in Table 4.
+For HIGHLIGHTS, highlight videos were generated using 300 traces, a context length of 5, and 20 highlights, while SySLLM summaries were generated from 50 captioned episodes (see Section 5).
 Condition Task 1 Sequence Task 1 Agent Type Task 2 Summary Type
@@
 4 HIGHLIGHTS $\rightarrow$ SySLLM Short-sighted Agent HIGHLIGHTS
-Table 6: Experimental conditions.
-7.3 Participants
-We recruited participants through Prolific (N = 200).
-Participants were native English speakers from the US, UK, Australia, or Canada.
-Participants received a 3.75 pound base payment and an additional, 1 pound bonus if participants answered all questions correctly in Task 2.
-Participants who failed the attention question were excluded from the final analysis, as well as participants who completed the survey in less time than the length of the videos presented.
-After exclusions, we had 192 participants (94 female, Mean age = 36.41, STD = 12.05).
-7.4 Results
-Task 1: Summary Preferences
-The SySLLM summaries consistently received higher ratings compared to the HIGHLIGHTS summaries across all metrics (see Figure 4).
-To assess the statistical significance of the differences in ratings, we first calculate the average scores for each participant under both the SySLLM and HIGHLIGHTS conditions.
-We then compute the difference between these averages for each participant.
-A paired t-test is conducted on these differences to evaluate whether the SySLLM condition significantly outperforms the HIGHLIGHTS condition across all metrics.
-The results indicate a statistically significant difference in favor of the SySLLM condition, with statistic $T=13.99$ and $p<10^{-33}$.
-Following the evaluations of specific summaries, we asked the participants to subjectively compare the textual and video summaries.
-Specifically, we asked them which summary they think is better for understanding the agent’s overall behavior based on the observed video of a large sample of trajectories.
-They were further asked to rate the extent to which they preferred the textual (or visual) summary (on a 1-7 Likert scale).
-The results we obtained are conclusive.
-The responses indicate a clear preference for the SySLLM summary, with 75.5% of participants favoring the language-based summaries.
-Regarding the second question, the mean rating was 5.97 with a standard deviation of 1.44.
-In the 1-7 Likert scale, a score of 4 represents indifference to the different policies; hence, a mean score of 5.97 indicates a substantial preference for the textual summary.
-Qualitative Analysis The open text feedback we received from participants of Task 1 turned out to be highly valuable.
-Some participants explained that while HIGHLIGHTS summaries required their own interpretation of the videos, SySLLM summaries provided an interpretation of the agent’s policy, e.g., “the [SySLLM] summary was thorough in explaining the movement of the agent.
-It describes information that isn’t visible in the video.
-The information in the summary was objective while watching the video summary and interpreting it is based on my view which is subjective.” Other participants mentioned that SySLLM summaries also explained why the agent acted as it did.
-For instance, one participant noted that “there are instances in the video where the agent seems to turn random corners instead of straight as you would expect it to.
-The summary explains why.” Another participant wrote that “the video itself does not fully explain but rather how you interpret what it is doing and in the case of errors it starts doing something wrong you don’t know what causes it”.
-Additionally, some participants liked that SySLLM summaries provided more context, e.g., “the textual summary provides context for what is occurring.
-Conversely, the video summary doesn’t inform the viewer of the ’behind the scenes’ that is occurring.
-From the textual information, I identified many strategies and code that the agent is using such as using the maze strategy to always turn right when it hits a wall.
-On the other hand, the video does not provide this information unless you have prior knowledge such as coding experience on conditionals and iterations to know what it occurring.”
-Figure 4: Participant scores for Task 1. Each metric includes a box plot of the Likert scores. Notably, SySLLM receives higher ratings than HIGHLIGHTS across all metrics.
-Task 2: Agent Identification
-As shown in Figure 5, participants’ ability to correctly identify the agent based on the provided summaries varied between the SySLLM and HIGHLIGHTS conditions across the three questions.
-For $Q1$ ($\chi^{2}=1.0523$, $p=0.3050$, odds ratio (OR) = 0.5843), $Q2$ ($\chi^{2}=0.6703$, $p=0.4129$, OR = 0.7550), and $Q3$ ($\chi^{2}=0.7279$, $p=0.3936$, OR = 1.3875), the Chi-Square tests showed no statistically significant differences.
-Independent t-tests showed no statistically significant differences in confidence scores between the SySLLM and HIGHLIGHTS conditions for $Q1$ (t=0.6202t=-0.6202, $p=0.5359$) and $Q2$ ($t=0.5565$, $p=0.5785$).
-However, for $Q3$, there was a statistically significant difference ($t=3.4197$, $p=0.0008$).
-To conclude, participants did not show a significant preference for the SySLLM or HIGHLIGHTS condition when identifying the correct agent based on the summaries provided, except for $Q3$ where the SySLLM condition had significantly higher confidence scores.
-Further investigation may be required to understand the nuances behind these results and whether different summary methods or conditions might yield different outcomes.
-Figure 5: Correctness rate in Task 2.
-For each question, we measure the proportion of users that identified the correct agent under both SySLLM, HIGHLIGHTS, and a random guess.
-For each bar, we included the 95% confidence intervals.
-Notice that both SySLLM and HIGHLIGHTS receive higher scores than the random guess.
-For all three questions, the difference in correctness rate between SySLLM and HIGHLIGHTS is not statistically significant.
-8 Discussion and Future Work
-In this paper, we introduced SySLLM, an innovative framework leveraging large language models to synthesize reinforcement learning policies into clear, structured narratives.
-Our method addresses key shortcomings of traditional extractive-visual methods by moving away from fragmented, scenario-specific snapshots toward holistic, language-based summaries.
-This shift significantly reduces the effort required for users to interpret complex agent behaviors, facilitating better comprehension of strategic decisions and adaptability.
-Our evaluations provide strong evidence that SySLLM effectively captures nuanced policy behaviors and aligns closely with expert-generated summaries.
-The qualitative alignment with human experts highlights the capability of language models to accurately convey intricate behavioral patterns and strategic nuances, reinforcing the effectiveness of a textual-abstractive approach.
-Furthermore, user preference studies demonstrate clear advantages in interpretability and user engagement over visual-based summarization methods.
-The broader implications of SySLLM suggest promising new directions for research and practice.
-We advocate for increased exploration into abstractive-textual explanations in global explanation in reinforcement learning, encouraging the community to investigate advanced techniques such as fine-tuning, in-context learning, and integration with reasoning frameworks.
-Embracing this paradigm can lead to significant advancements in how we communicate complex policies, ultimately enhancing trust, transparency, and adoption of reinforcement learning in critical real-world scenarios.
-Captioners
-Our approach relies on the availability of effective captioning functions, a common requirement in language-integrated methodologies as seen in prior research [35, 46, 45].
-In simulated environments, this dependency is relatively straightforward to address, as ground truth states are accessible, and many RL environments—such as NetHack/MiniHack [56, 57] and text-based games [58, 59]—already incorporate language-based features.
-For more complex or real-world scenarios, however, implementations would require robust object detection [60], advanced captioning frameworks [61], action recognition systems [62], or vision-language models [63].
-Alternatively, annotator models [35] or pretrained foundation models [64] could generate captions.
-The continued advancement of general-purpose captioning models suggests that off-the-shelf solutions will soon support an even broader range of tasks, enhancing the practicality of our approach.
-LLMs Limitations
-The current performance of SySLLM is bounded by the capabilities and limitations of the underlying language models.
-Context window constraints restrict the amount of agent trajectory data that can be analyzed at once, potentially limiting the ability to identify long-term patterns or rare behaviors.
-Additionally, LLMs can occasionally generate plausible-sounding but incorrect explanations (hallucinations), particularly when faced with unusual or complex agent behaviors.
-As LLM technology evolves, particularly with expanded context windows and improved factuality, we anticipate that these limitations will diminish, enabling more comprehensive and reliable agent behavior summaries.
-Building on SySLLM’s strengths while addressing its limitations, we identify several promising directions for future research:
-Interactive Question-Answering Framework
-A promising avenue for future work is extending SySLLM to support an interactive question-answering (QA) framework.
-This would enable users to directly query the agent’s behavior, asking targeted questions such as “Why did the agent choose action A over B in this situation?” or “How would the agent respond to condition X?” Such interactivity could transform SySLLM from a static summarization tool into a dynamic interface for policy exploration, offering a more flexible and intuitive means of understanding agent behavior.
-Prior work on logical querying of agents, such as [65], highlights the feasibility and potential of QA-based explanations in RL.
-Inspired by these developments, we envision a natural language-based QA interface grounded in narrative trajectories, allowing users to probe specific aspects of agent decision-making in a more accessible way.
-Multi-modal Summaries
-Integrating natural language explanations with visual representations would significantly enhance the communicative power of SySLLM.
-For each behavioral pattern identified by the language model, the system could provide corresponding visual evidence extracted from agent traces such as state transitions, action sequences, or critical decision points.
-This multi-modal approach would ground abstract explanations in concrete examples, potentially mitigating hallucination issues while making complex policies more accessible to users with diverse backgrounds and expertise levels.
-Comparative Analysis Framework
-Future work could extend SySLLM to automatically compare multiple agents’ behaviors, identifying similarities, differences, and relative strengths.
-This comparative capability would be particularly valuable for evaluating alternative training approaches, understanding the effects of hyperparameter changes, or analyzing how agent behavior evolves during training.
-By articulating these comparisons in natural language, SySLLM could provide intuitive insights into the factors that influence agent performance and behavior.
-Acknowledgments
-This project was funded by the ERC Starting Grant #101078158 CONVEY and the Israel Science Foundation (ISF) under Grant No. 3079/24.
+Table 4. Experimental conditions.
+Procedure.
+The participants first completed a tutorial on MiniGrid Unlock rules, followed by a comprehension quiz.
+In Task 1, they watched a 120 seconds video of an agent’s behavior and then rated a summary (SySLLM or HIGHLIGHTS) on eight explanation quality metrics (7-point Likert), adapted from Hoffman et al.
+hoffman2018metrics.
+After evaluating both modalities, they provided direct preference judgments: which summary better reflected the agent’s policy and by what margin.
+In Task 2, participants were shown a summary (textual or visual) and asked to match it to one of three short (20s) videos: the correct agent plus two distractors.
+Each participant completed three trials ($Q1$: goal-directed, $Q2$: random, $Q3$: short-sighted).
+For each, they indicated their choice, rated confidence, and provided a justification.
+Participants.
+We recruited 200 participants from Prolific (native English speakers from the US, UK, Canada, and Australia).
+Compensation was £3.75 base plus a £1 bonus for correct completion of Task 2.
+After exclusions for failed attention checks and implausibly short completion times (below 300 seconds), 192 participants remained (94 female, $M_{age}=36.4$, $SD=12.1$).
+Results.
+In Task 1, SySLLM consistently outperformed HIGHLIGHTS across all metrics (Fig.
+5).
+Paired $t$-tests confirmed the difference as highly significant ($T=13.99$, $p<10^{-33}$).
+Direct preference questions reinforced this: 75.5% of participants favored SySLLM, and the comparative Likert rating averaged $M=5.97$, $SD=1.44$ (neutral baseline = 4).
+Qualitative feedback highlighted that SySLLM explained why agents acted as they did, while HIGHLIGHTS required subjective inference.
+For example: “There are instances in the video where the agent seems to turn random corners.
+The summary explains why.” In Task 2, correctness rates for both modalities exceeded random-guess baselines (Fig.
+6).
+Chi-Square tests found no significant differences between SySLLM and HIGHLIGHTS across Q1–Q3.
+However, confidence scores revealed a significant effect in Q3, where participants in the SySLLM condition reported higher confidence ($t=3.42$, $p=0.0008$).
+Overall, participants rated SySLLM summaries as significantly clearer and more informative than the highlight videos.
+Although both modalities supported correct agent identification, textual summaries provided stronger interpretive cues, particularly reflected in higher confidence for certain agents.
+These results suggest that abstractive, language-based policy summaries enhance subjective interpretability while maintaining competitive performance in behavior recognition.
+[…]
+Participant scores for Task 1.
+Figure 5. Participant ratings for Task 1 on a 1–7 Likert scale. SySLLM ratings are significantly higher than HIGHLIGHTS ratings across all metrics.
+[…]
+Correctness rate in Task 2.
+Figure 6.
+Correctness rate in Task 2.
+Error bars indicate 95% confidence intervals.
+Both SySLLM and HIGHLIGHTS outperform the random guess baseline.
+The differences in correctness rate between SySLLM and HIGHLIGHTS is not statistically significant.
+9. Discussion and Future Work
+We introduced SySLLM, a framework for abstractive textual explanation of reinforcement learning policies.
+By converting state and action trajectories into structured language and leveraging large language models to generate policy summaries, SySLLM abstracts low-level decision traces into high-level behavioral patterns.
+Empirical results show close alignment with expert interpretations and a clear user preference over visual demonstrations, supporting language-based abstraction for interpreting complex policies.
+Several limitations remain.
+SySLLM currently relies on domain-specific captioning functions to translate trajectories into text.
+While feasible in simulated environments, extending to high-dimensional or partially observed domains will require perceptual grounding via visual-language pipelines or pretrained vision-language models capable of zero-shot scene understanding.
+Future work extends beyond static summarization.
+One direction is interactive policy querying, where users ask natural language questions about agent behavior, enabling counterfactual and rationale-based explanations.
+Another is comparative summarization, contrasting policies across training stages, reward functions, or architectures.
+Finally, integrating textual and visual modalities and advancing toward open-domain policy summarization may enable benchmark-agnostic systems that characterize agent behavior at scale.
+{acks}
+Funded by the European Union (ERC, Convey, 101078158) and the Israel Science Foundation (ISF) under Grant No.
+3079/24.
+Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency.
+Neither the European Union nor the granting authority can be held responsible for them.
 Appendix A HIGHLIGHTS
 “Highlights” Policy Summaries
-Our user study uses “Highlights” policy summaries [17] as a baseline.
+Our user study uses “Highlights” policy summaries amir2018highlights as a baseline.
 For completeness, we describe the algorithm here.
@@
 Cumulative Reward The total reward accumulated by the agent up to that step.
-Table 7: Description of the data stored in the experience dataset.
+Table 5. Description of the data stored in the experience dataset.
 Appendix C Summarization Prompt:
@@
 Appendix F Implementation Details
-F.1 MiniGrid
+F.1. MiniGrid
 We employed the PPO algorithm from the stable-baselines3 library for our policy network, which takes as input a $K\times K\times 3$ encoded image and a mission string, the latter being encoded using a one-hot scheme.
@@
 Clip Range 0.2 0.2 0.2 0.2 0.2 0.2
-Table 8: Hyper-parameters for the PPO algorithm applied to all six agents.
-F.2 Crafter
+Table 6. Hyper-parameters for the PPO algorithm applied to all six agents.
+F.2. Crafter
 We implemented DreamerV3 for our agent, using a state-of-the-art world model-based reinforcement learning approach.
@@
 The questions were phrased according to the condition—either video or natural language summary.
-  1. 1.
+  1. (1)
     Clarity: “The [video/natural language] summary clearly explained the agent’s actions and decisions shown in the demonstration video.”
-  2. 2.
+  2. (2)
     Understandable: “From the [video/natural language] summary, I understand how the agent’s actions and decisions shown in the demonstration video.”
-  3. 3.
+  3. (3)
     Completeness: “The [video/natural language] summary seemed complete in covering all aspects of the agent’s actions and decisions in the demonstration video.”
-  4. 4.
+  4. (4)
     Satisfaction: “The [video/natural language] summary is satisfying in capturing the agent’s behavior and decisions displayed in the demonstration video.”
-  5. 5.
+  5. (5)
     Useful: “The [video/natural language] summary is useful to my understanding of the agent’s behavior and decisions displayed in the demonstration video.”
-  6. 6.
+  6. (6)
     Accuracy: “The information in the [video/natural language] summary accurately reflected the agent’s behavior and decisions displayed in the demonstration video.”
-  7. 7.
+  7. (7)
     Improvement: “The [video/natural language] summary provides additional insights about the agent’s behavior that are not immediately apparent from watching the demonstration video alone.”
-  8. 8.
+  8. (8)
     Preference: “I prefer receiving information about agent behavior through the [video/natural language] summary rather than just watching the demonstration video.”
@@
 This process involved a quantitative evaluation of prompt designs based on observed outputs, guided by principles from prompt engineering literature, and tailored to domain-specific requirements.
-Additionally, the final design was inspired by the Chain of Thought (CoT) [47] prompting paradigm, which encourages models to generate structured, step-by-step reasoning.
+Additionally, the final design was inspired by the Chain of Thought (CoT) wei2022chain prompting paradigm, which encourages models to generate structured, step-by-step reasoning.
 Below is a detailed breakdown of the methodology used: